### v2

# Fenchel-Young Variational Inference

Abstract
Bayesian learning and inference seek distributions of the unknowns—parameters and/or latent variables—that balance empirical risk and regularization. In both standard and variational Bayesian inference, this disagreement is measured by the Kullback-Leibler divergence (KLD). We present Fenchel-Young variational inference (FYVI), which broadens previous formulations by replacing the KLD with a Fenchel-Young (FY) regularizer. We focus on a class of FY losses based on Tsallis entropies that contains the standard KLD as a particular case, and which allows us to obtain generalized variational posterior distributions with support strictly smaller than that of the prior. FYVI includes new generalized notions of free energy, evidence, and posterior. We derive alternating minimization and gradient backpropagation algorithms to approximate the generalized FY posterior. Our FY expectation-maximization algorithm and FY latent variable models can be seen as generalized variants of classical algorithms and models. The proposed methods are empirically competitive, outperforming established counterparts in controlled experiments, in which directly modeling limited support (sparse) posteriors is intuitively appealing.
1 Introduction
Much of machine learning can be formulated as variational optimization problems aiming to minimize a regularized empirical risk (average loss) over a training set, where the regularizer encourages the solution—learned parameters—to have certain properties. In Bayesian approaches to machine learning and inference, all the unknowns—learnable parameters and, possibly, latent/hidden variables—are seen as random variables, and learning becomes synonymous with inferring the conditional distribution of the unknowns given the observed data. Based on the variational formulation of Bayesian inference by Zellner [59] and the generalized loss posteriors of Bissiri et al. [7], Khan and Rue [32] unify diverse, well-established Bayesian inference algorithms under a common formal umbrella: finding the (generalized) posterior distribution (in some allowed family) of the unknown parameters or latent variables that strikes a balance between a low expected risk and a small Kullback-Leibler divergence (KLD) from a given prior.
In this paper, we extend the variational formulations of (generalized) posterior inference studied along related lines by Bissiri et al. [7], Knoblauch et al. [34], and Khan and Rue [32], by considering alternative optimality criteria that exploit the class of Fenchel-Young (FY) losses [10, 39]. In particular, we focus on FY losses based on Tsallis entropies, which subsume the Kullback-Leibler divergence as a particular case. We generalize the formulations of Knoblauch et al. [34] and Khan and Rue [32], leading to a strictly wider class of statistical families, which includes generalized posteriors with limited (non-full, a.k.a. sparse) support, as well as heavy-tailed posteriors. [^1] In this paper, we focus on the less studied case of sparse posteriors, rather than on heavy-tailed ones, which are a non-trivial case that also falls under the framework we propose.
As reviewed in the next section, Zellner [59] rediscovers the fact that the Bayesian posterior is the solution of an unconstrained optimization/variational problem over the space of distributions where the objective function(al) is the expected log-likelihood plus the Kullback-Leibler divergence (KLD) from the prior [20, 16]. The KLD term in the variational formulation of posterior inference is equivalent to a maximum entropy regularizer that favors posteriors with high Shannon entropy [31], giving a privileged position to exponential families in this context [25], following from the well-known fact that these are maximum Shannon entropy distributions, under expectation constraints.
If instead of the KLD, we use FY losses induced by Tsallis entropies, the corresponding generalized posteriors belong to deformed exponential families, which, unlike standard exponential families, may have variable support. Whereas Knoblauch et al. [34] analyze arbitrary combinations of divergences, losses, and exponential variational families, we show that when using FY losses based on Tsallis entropies, the variational family cannot be chosen independently of the divergence; in fact, it is determined by it.
In §3 we propose the Fenchel-Young variational infrence (FYVI) framework, which includes new generalized notions of FY free energy, FY evidence, FY evidence lower bound, and FY posterior. We derive an EM-style [19] algorithm for parameter estimation, and describe gradient backpropagation for amortized inference with deformed exponential family posteriors. Both of these estimation methods can be used to fit arbitrary generative models that may be specified as instances of FYVI.
In §4 we present experiments with the FYEM algorithm applied to Gaussian mixture model (GMM) [42] estimation, and two FY variational autoencoder models (VAEs) [33, 50], one applied to images and the other to documents. The FYEM algorithm for GMMs recovers the classical EM as a special case and yields a new generalization with an adaptively sparse E-step. In the image FYVAE we obtain the best performance when we combine a sparse posterior with a sparse observation model. We also outperform the VAE baseline of a neural variational document model (NVDM) [40], using a FY loss in place of a negative log-likelihood combined with a standard Gaussian latent variable, by tuning the sparsity of the observation model.
Our code will be publicly released upon acceptance.
2 Background
2.1 Definitions and Notation
[…]
2.2 Variational View of Bayesian Inference
2.2.1 Standard Bayesian Inference
As has been shown [12, 16, 20, 59], the Bayesian posterior $p_{Z|X}$ can be expressed in variational form as the distribution in $\Delta(\mathcal{Z})$ which strikes a balance between having a small KLD from the prior $p_{Z}$, and yielding a high expected log-likelihood $\log p_{X|Z}$, given some observation $x\in\mathcal{X}$. Formally,
${p_{Z|X}(\cdot|x)}$ $=\arg\!\!\max_{q\in\Delta(\mathcal{Z})}\underset{Z\sim q}{\mathbb{E}}\bigl{[}\log p_{X|Z}(x|Z)\bigr{]}\!-\!D_{\textnormal{KL}}(q\|p_{Z})$ (1)
$=\arg\!\!\max_{q\in\Delta(\mathcal{Z})}\underset{Z\sim q}{\mathbb{E}}\bigl{[}\log p_{X,Z}(x,Z)\bigr{]}\!+\!H(q).$ (2)
The optimum of (1), achieved by setting $q=p_{Z|X}(\cdot|x)$, is (after trivial manipulation) equal to the log-evidence:
$\mathbb{E}_{Z\sim p_{Z|X}(\cdot|x)}\bigl{[}\log p_{X,Z}(x,Z)\bigr{]}+H\bigl{(}p_{Z|X}(\cdot|x)\bigr{)}=\log p_{X}(x).$ (3)
More often than not, the posterior and evidence cannot be easily obtained, which has stimulated decades of work on approximate Bayesian methods.
2.2.2 Variational Inference
In variational inference (VI) [9, 31], the goal is to approximate the posterior and bound the evidence when these are intractable, by confining $q$ to a tractable family, i.e. a subset of the entire space of probability functions, $\mathcal{Q}\subseteq\Delta(\mathcal{Z})$. The familiar VI optimization problem is
$q_{x}^{\star}=\arg\max_{q\in\mathcal{Q}}\underset{Z\sim q}{\mathbb{E}}\bigl{[}\log p_{X,Z}(x,Z)\bigr{]}+H(q),$ (4)
and its optimal value is the ELBO (evidence lower bound): ${\mathbb{E}_{Z\sim q_{x}^{\star}}}\bigl{[}\log p_{X,Z}(x,Z)\bigr{]}+H\bigl{(}q_{x}^{\star}\bigr{)}\leq\log p_{X}(x)$. The negative of the objective function in (4) is often called the variational free energy.
2.2.3 Generalized Variational Inference
Khan and Rue [32] use Bissiri et al. [7]’s extension of the classical variational problem in (4) to generalized posteriors, in order to derive as special cases much of the core catalog of modern machine learning algorithms and to propose new ones. This extension consists of replacing the negative log-likelihood with an arbitrary loss function $\ell(x;Z)$:
$q_{x}^{\star}=\arg\min_{q\in\mathcal{Q}}\underset{Z\sim q}{\mathbb{E}}\left[\ell(x;Z)\right]+D_{\text{\footnotesize KL}}(q\|p_{Z})=\arg\min_{q\in\mathcal{Q}}\underset{Z\sim q}{\mathbb{E}}\left[\ell(x;Z)-\log p_{Z}(Z)\right]-H(q),$ (5)
where the notation $q_{x}^{\star}$ highlights the dependency on the observations $x$. Khan and Rue [32] restrict $\mathcal{Q}$ to be a regular minimal exponential family, i.e., $\mathcal{Q}=\{q(z)=h(z)\exp\bigl{(}\langle\lambda,t(z)\rangle-A(\lambda)\bigr{)},\mbox{for }\lambda\in\Lambda\}$, where $\Lambda$ is the parameter space, $t$ is a canonical sufficient statistic, $h$ is the base measure (independent of $\lambda$), and $A(\lambda)$ is the log-partition function. In a similar vein, Knoblauch et al. [34] propose generalized VI (GVI), by additionally considering divergences other than KLD, and introducing the rule of three (RoT), which refers to the choice of the three elements required for GVI: a loss function, a divergence measuring deviation of variational posterior from prior, and a set of allowed solutions $\mathcal{Q}$.
2.3 Fenchel-Young Losses and Related Work
In a previously unrelated line of research, Blondel et al. [10] propose Fenchel-Young (FY) losses as a generalization of the ubiquitous cross-entropy loss that can be seen as mixed space variants of Bregman divergences [4], [10, §3.2]. FY losses have been used for supervised classification problems, but not in any kind of (generalized) Bayesian contexts, such as VI or latent variable models. Our notation below treats both discrete and continuous domains in a unified way.
Let $\Omega:\Delta(\mathcal{Z})\rightarrow\bar{\mathbb{R}}$ be a lower semi-continuous (l.s.c.) proper convex function, and let $\Omega^{*}$ be its Fenchel-Legendre conjugate [^2] Given a function $f:\mathcal{U}\rightarrow\bar{\mathbb{R}}$, its Fenchel conjugate $f^{*}:\mathcal{U}^{*}\rightarrow\bar{\mathbb{R}}$ is defined as $f^{*}(x)=\sup_{u\in\mathcal{U}}(\langle x,u\rangle-f(u)),$ where $\mathcal{U}^{*}$ is the dual space of $\mathcal{U}$. If $f$ is convex l.s.c., biconjugation holds: $f^{**}=f$. If $\mathcal{U}$ is compact or $f$ is continuous, sup may be replaced with max, because the supremum is achieved. In this case, assuming differentiability, the maximizing argument in the definition of $f^{*}$ is $\arg\max_{u\in\mathcal{U}}(\langle x,u\rangle-f(u))=\nabla f^{*}(x)$. [11, 6]. Consider also a score function $\eta:\mathcal{Z}\rightarrow\mathbb{R}$ that assigns a real-valued score to each element of $\mathcal{Z}$. In a Bayesian setting, $\eta$ corresponds to a log-prior. For a distribution $q\in\Delta(\mathcal{Z})$, the FY loss induced by $\Omega$ is a measure of how incompatible $\eta$ and $q$ are:
$L_{\Omega}(\eta;q)=\Omega^{*}(\eta)-\langle q,\eta\rangle+\Omega(q)=\Omega^{*}(\eta)-\mathbb{E}_{Z\sim q}[\eta(Z)]+\Omega(q),$ (6)
where the second equality follows from the simple fact that $\mathbb{E}_{Z\sim q}[\eta(Z)]=\langle q,\eta\rangle$, for discrete or continuous $\mathcal{Z}$. An FY loss is always non-negative, due to the Fenchel-Young inequality [6],
$\Omega^{*}(\eta)=\max_{q\in\Delta(\mathcal{Z})}\bigl{(}\langle q,\eta\rangle-\Omega(q)\bigr{)}\;\Rightarrow\;\Omega^{*}(\eta)-\langle q,\eta\rangle+\Omega(q)\geq 0,$
where max can be used instead of sup, because $\Delta(\mathcal{Z})$ is compact. Furthermore, $L_{\Omega}(\eta;q)$ is zero if and only if $q=\arg\max_{q\in\Delta(\mathcal{Z})}\mathbb{E}_{Z\sim q}[\eta(Z)]-\Omega(q)$. This maximization defines the $\Omega$-regularized prediction map $\Pi_{\Omega}:\mathcal{F}\rightarrow\Delta(\mathcal{Z})$,
$\Pi_{\Omega}(\eta):=\arg\max_{q\in\Delta(\mathcal{Z})}\mathbb{E}_{Z\sim q}[\eta(Z)]-\Omega(q),$ (7)
where $\mathcal{F}\subseteq\mathbb{R}^{\mathcal{Z}}$ is the set of functions for which the maximizer exists and is unique [10]. A trivial but important property of $\Pi_{\Omega}$ is invariance under global shifts of the argument: $\Pi_{\Omega}[\eta+C]=\Pi_{\Omega}[\eta]$, for any constant $C.$
[…]
A particularly relevant choice for $\Omega$, which we will adopt in this paper, is the Tsallis negentropy [53],
$\Omega_{\rho}(q):=\left\{\begin{array}[]{ll}\displaystyle\mathbb{E}_{Z\sim q}[q(Z)^{\rho-1}-1]/\bigl{(}\rho(\rho-1)\bigr{)},&\text{if $\rho\neq 1$,}\\ \mathbb{E}_{Z\sim q}[\log q(Z)],&\text{if $\rho=1$,}\end{array}\right$. (10)
where $\rho>0$ is a parameter, called “entropic index”. Tsallis negentropies are continuous on $\rho$ and coincide with the Shannon negentropy when $\rho=1$. The $\Omega_{\rho}$-regularized prediction map $\Pi_{\Omega_{\rho}}$ is called $\rho$-entmax [47] and recovers softmax and sparsemax, for $\rho=1$ and $\rho=2$, respectively. For any $\rho>1$, $\Omega_{\rho}$-regularized prediction map will be sparse, with higher values of $\rho$ giving distributions with increasingly sparse supports.
Appendix A describes the relationship between FY losses based on Tsallis negentropies and the $\alpha$- and $q$-divergences of Amari and Ohara [4]. Amari’s $\alpha$-divergences, which have connections with Tsallis statistics, are related to but distinct from the divergences associated with Rényi entropies, which, somewhat confusingly, are also called $\alpha$-divergences; we do not address Variational Rényi approaches in this paper [36, 41, 34, 17, 18, 26].
3 Fenchel-Young Variational Inference
3.1 Fenchel-Young Variational Posteriors
Our proposed FYVI formulation is built by substituting the Kullback-Leibler divergence in (5) with the FY loss induced by a l.s.c. regularizer $\Omega$ with respect to a log-prior $\eta=\log p_{Z}$. This yields a new class of generalized posterior distributions of the following form:
$q_{x}^{\star}:=\arg\min_{q\in\mathcal{Q}}\underset{Z\sim q}{\mathbb{E}}\left[\ell(x;Z)\right]+L_{\Omega}(\eta;q)=\arg\min_{q\in\mathcal{Q}}\underset{Z\sim q}{\mathbb{E}}\left[\ell(x;Z)-\eta(Z)\right]+\Omega(q).$ (11)
At this point, we do not restrict or specify $\mathcal{Q}$.
In this paper, we focus on FYVI with Tsallis $\rho$-negentropies $\Omega_{\rho}$ as regularizers, which as mentioned above, for $\rho>1$, naturally induce variational posteriors with adaptively sparse support, i.e., such that the support may be strictly smaller than $\mathcal{Z}$ and may depend on $x$. In contrast, for $\rho<1$, the variational posteriors may be denser (closer to uniform) than those obtained with standard Shannon entropy regularization (i.e., $\rho=1$); we will not consider this case, and leave it for future work.
By definition, an exponential family has a support that cannot depend on its parameters [5]. Consequently, due to their variable support, generalized posteriors obtained via (11), where $\Omega=\Omega_{\rho}$, a Tsallis $\rho$-negentropy with $\rho>1$, cannot belong to an exponential family. This fact implies that contrary to the suggestion of the RoT (rule of three [34]), the variational family $\mathcal{Q}$ and the divergence cannot always be chosen independently of each other. Namely, adopting a divergence that may yield generalized posteriors with variable/adaptive support is incompatible with a variational family in any exponential family. In Appendix A, following similar derivations as Martins et al. [39, Prop. 8], we prove Proposition 1, demonstrating that for $\Omega=\Omega_{\rho}$, problem (11) is equivalent to a maximum Tsallis maximum $\rho$-entropy problem, the solution of which belongs to a $(2-\rho)$-exponential family,
$q_{x}^{\star}(z)$ $=\left[1+(1-\rho)\left(\eta(z)-\ell(x;z)-A_{\rho}\right)\right]_{+}^{\frac{1}{\rho-1}}=\exp_{2-\rho}\left[\eta(z)-\ell(x;z)-A_{\rho}\right],$ (12)
where $A_{\rho}$ is a normalization function and $a]_{+}=\max\{0,a\}$ is the positive part operator, a.k.a. rectified linear unit (ReLU). Notice that the ReLU operation may produce zero results, thus assigning exactly zero probability (or probability density) to some elements of $\mathcal{Z}$, and the support of $q_{x}^{\star}$ may depend on $x$, something that is impossible in any exponential family.
We stress that FYVI, with Tsallis $\rho$-negentropy regularization ($\rho>1$), leads to a qualitatively different type of sparsity from what is obtained by using sparsity-inducing priors or regularizers, such as LASSO—equivalently, Laplacian priors [52, 27]—or spike-and-slab priors [30]. Sparsity-inducing priors are used to encourage the estimated object—typically a vector of parameters of coefficients—to be sparse, but the corresponding posterior distributions have full support. For example, the Bayes posterior resulting from a Laplacian sparsity-inducing prior and a Gaussian observation model has full support. The reason why it is said to induce sparsity is that the posterior mode (the maximum a posteriori (MAP) estimate) can be exactly zero. In contrast, even using a non-sparsity-inducing Gaussian prior and a Gaussian observation model, together with FYVI with, say, Tsallis $2$-negentropy regularization, yields a generalized posterior density with bounded support, since it results from the application of the 2-entmax (sparsemax) prediction map [39].
3.2 Fenchel-Young Free Energy, Evidence, Evidence Lower Bound, and Posterior
By analogy with (4) the objective in (11) can be seen as a Fenchel-Young variational free energy (FYVFE). It can be useful to include an additional parameter $\beta$ weighting the KLD regularizer in (5) or the FY loss below, as in the $\beta$-variational auto-encoders ($\beta$-VAE [28]):
$F_{\ell,\Omega,\beta}(q,\eta;x)=\underset{Z\sim q}{\mathbb{E}}\left[\ell(x;Z)\right]+\beta L_{\Omega}(\eta;q).$ (13)
The parameter $\beta$ controls the strength with which the agreement between $q$ and $\eta$ is encouraged, and can also be seen from an information-theoretic perspective as controlling a rate-distortion trade-off [2]. In the sequel, omitting $\beta$ means assuming $\beta=1$.
By analogy with (3), the minimum of 13 is the (negative) FY evidence,
$J_{\ell,\Omega}(x,\eta)=\min_{q\in\mathcal{Q}}F_{\ell,\Omega}(q,\eta;x)=F_{\ell,\Omega}(q_{x}^{\star},\eta;x),$
and the minimizer $q_{x}^{\star}$ is the FY posterior. Since, by construction, the FY evidence $J_{\ell,\Omega}(x,\eta)$ is the maximum achieved by $F_{\ell,\Omega}(q,\eta;x)$, for any $q\in\Delta(\mathcal{Z}),$ the negative free energy $F_{\ell,\Omega}(q,\eta;x)$ is a FY evidence lower bound (FYELBO). For $\mathcal{Q}=\Delta(\mathcal{Z})$, $\Omega(q)=\Omega_{1}(q)=-H(q)$, and $\ell(x;z)=-\log p_{X|Z}(x|z)$, the FY evidence becomes the regular log-evidence $\log p_{X}(x)$ and the FY posterior recovers the standard Bayesian posterior $q_{x}^{\star}=p_{Z|X}(\cdot|x)$.
Although the ELBO (standard or FY) designation suggests that VI aims at bounding the evidence, the main goal of VI (standard or FY) is not to bound or approximate the evidence but rather to find the best possible approximation to the posterior in some variational family. In classical Bayesian analysis, a prior on the model space allows us to use the evidence for model comparison or hypothesis testing [37]. However, the FY evidence lacks the standard statistical interpretation as a marginal likelihood, and Bayesian model selection, validation, and hypothesis testing are not addressed here.
3.3 Fenchel-Young Expectation Maximization (FYEM)
Suppose $x$ contains $N$ independent observations, $x=(x_{1},...,x_{N})\in\mathcal{X}^{N}$, associated with a collection of latent variables $z=(z_{1},...,z_{N})\in\mathcal{Z}^{N}$, with corresponding variational distributions $q=(q_{1},...,q_{N})\in\Delta(\mathcal{Z})^{N}$. Let $\ell_{\theta}(x;z)=\sum_{i=1}^{N}\mathcal{L}_{\theta}(x_{i};z_{i})$ be a parameterized loss function, which is the sum of sample-wise losses, where $\theta$ is a vector of parameters to be estimated. The FYVI problem is to minimize the FYVFE (13) with respect to $q$ and $\theta$,
$\hat{q},\hat{\theta},\hat{\eta}$ $=\arg\min_{q,\theta,\eta}\sum_{i=1}^{N}F_{\ell_{\theta},\Omega}(q_{i},\eta;x_{i})=\arg\min_{q,\theta,\eta}\sum_{i=1}^{N}\Bigl{(}\underset{Z_{i}\sim q_{i}}{\mathbb{E}}\left[\mathcal{L}_{\theta}(x_{i};Z_{i})\right]+L_{\Omega}(\eta;q_{i})\Bigr{)}.$ (14)
where $\eta$ is a score function (possibly a log-prior), also to be estimated.
Alternating minimization w.r.t. $q$ and $(\theta,\eta)$ yields an EM-style algorithm: the E-step computes the posteriors of the latent variables; as in standard EM, each $q_{i}$ can be obtained independently of the others. The M-step updates the estimates of $\theta$ and $\eta$.
E-step:
Given the current $\hat{\theta}$ and $\hat{\eta}$, update $\hat{q}_{i}$ by minimizing (14),
$\hat{q}_{i}^{\mbox{\scriptsize new}}$ $=\arg\min_{q_{i}}\underset{Z_{i}\sim q_{i}}{\mathbb{E}}\left[\mathcal{L}_{\hat{\theta}}(x_{i};Z_{i})\right]+L_{\Omega}(\hat{\eta};q_{i})=\arg\max_{q_{i}}\underset{Z_{i}\sim q_{i}}{\mathbb{E}}\left[\hat{\eta}(Z_{i})-\mathcal{L}_{\hat{\theta}}(x_{i};Z_{i})\right]-\Omega(q_{i})$
$=\arg\min_{q_{i}}L_{\Omega}(\hat{\eta}-\mathcal{L}_{\hat{\theta}}(x_{i};\cdot);q)=\Pi_{\Omega}(\hat{\eta}-\mathcal{L}_{\hat{\theta}}(x_{i};\cdot)),$ (15)
for i=1,,Ni=1,...,N, where $\Pi_{\Omega}$ is the $\Omega$-regularized prediction map (7), namely softmax, sparsemax, or another $\rho$-entmax transformation.
M-Step:
The M-step corresponds to updating $\hat{\theta}$ and $\hat{\eta}$, given $\hat{q}_{i}^{\mbox{\scriptsize new}}$. To update $\hat{\theta}$, we solve
$\hat{\theta}^{\mbox{\scriptsize new}}=\arg\min_{\theta}\sum_{i=1}^{N}\underset{Z_{i}\sim\hat{q}_{i}^{\mbox{\scriptsize new}}}{\mathbb{E}}[\mathcal{L}_{\hat{\theta}}(x_{i};Z_{i})],$
which depends on the particular form of $\mathcal{L}_{\hat{\theta}}$. To update $\hat{\eta}$, recall (see (6)) that $L_{\Omega}(\eta;q)=\Omega^{*}(\eta)-\mathbb{E}_{Z\sim q}[\eta(Z)]+\Omega(q)$, thus
$\hat{\eta}^{\mbox{\scriptsize new}}$ $=\arg\min_{\eta}\sum_{i=1}^{N}L_{\Omega}\bigl{(}\eta;\hat{q}_{i}^{\mbox{\scriptsize new}}\bigr{)}=\arg\min_{\eta}\Omega^{*}(\eta)-\frac{1}{N}\sum_{i=1}^{N}\underset{Z_{i}\sim\hat{q}_{i}^{\mbox{\scriptsize new}}}{\mathbb{E}}[\eta(Z_{i})]$
$=\arg\max_{\eta}\underset{Z\sim\bar{q}}{\mathbb{E}}[\eta(Z)]-\Omega^{*}(\eta),$ (16)
where $\bar{q}=(1/N)\sum_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}$, since the average of expectations under a set of probability (mass or density) functions is the expectation under the average of these functions. Assuming $\mathcal{Z}=\{1,...,K\}$ is a discrete set, $\eta$ is simply a vector $\eta=(\eta_{1},...,\eta_{K})\in\mathbb{R}^{K}$ and $q_{i}=(q_{i1},...,q_{iK})\in\Delta(\mathcal{Z})$ is probability distribution. In this case, $\mathbb{E}_{Z\sim\bar{q}}[\eta(Z)]=\langle\bar{q},\eta\rangle=\sum_{k=1}^{K}\bar{q}_{k}\eta_{k}$, thus
$\hat{\eta}^{\mbox{\scriptsize new}}\in\arg\min_{\eta}\;\Omega^{*}(\eta)-\langle\bar{q},\eta\rangle,$ (17)
where we have written the update as an inclusion because, as shown by Martins et al. [39], the minimizer may not be unique and is achieved by any $\hat{\eta}$ satisfying $\Pi_{\Omega}(\hat{\eta})=\bar{q}$. If $\Omega$ is the Shannon negentropy, $\Pi_{\Omega}$ is softmax, and $\hat{\eta}^{\mbox{\scriptsize new}}=(\log\bar{q}_{1},...,\log\bar{q}_{K})+C$, where $C$ is an arbitrary constant, which, without loss of generality, we set to zero, since it does not affect the algorithm. If $\Omega$ is the Tsallis 2-negentropy $\Omega_{2}$, $\Pi_{\Omega}(\eta)$ is the Euclidean projection of $\eta$ onto the simplex, and any $\hat{\eta}$ of the form $\hat{\eta}^{\mbox{\scriptsize new}}=(\bar{q}_{1},...,\bar{q}_{K})+C$, where $C$ is again an arbitrary constant that we set to zero, is a solution (although there may exist other solutions).
3.4 Gradient Backpropagation
In two of the experiments discussed below, we optimize the FYELBO in an amortized inference setting [24], where the variational distribution is parametrized by an encoder network that takes $x$ as input, as in variational autoencoders (VAE) [33]. We denote the resulting variational distribution by $q_{\phi}(z|x)$, where $\phi$ are the inference network parameters. Instead of minimizing (14) through alternating minimization, we use end-to-end gradient backpropagation combined with a reparametrization trick to update the encoder parameters $\phi$, the decoder parameters $\theta$, and, optionally, any parameters associated with $\eta$ (our FYVAEs use standard fixed priors). It is also obviously possible to use a score function estimator instead of reparametrization.
4 Experiments
We illustrate FYVI through in several applications. The first is GMM-based clustering, where we recover the standard and hard versions of EM for particular choices of regularizer, in addition to a new sparse [^4] This sparse variant of EM is a different instance of FYVI than those proposed by Dukkipati et al. [21] and Inoue and Shinoda [29], which are standard mixtures of $(2-\rho)$-Gaussians. EM. We then use FYVI in $\beta$-VAEs with finite posterior support and apply them to MNIST image data with both sparse and standard multivariate Bernoulli observation models, and to Fashion MNIST data with sparse and standard Categorical observation models. Finally, we combine a Gaussian prior with sparse observation models to model a sparse empirical distribution in a neural variational document model.
Our focus is not computational scalability. Depending on the particular model, parameter $\rho$ in Tsallis $\rho$-negentropy, and parametrization, the computational costs will be different. Computational cost depends on model-specific choices, rather than the generalization of standard VI to FYVI. Tradeoffs between FY and KL regularization similarly depend on model specification; there is no one general comparison. In most of our experiments, computational tradeoffs between FY regularizers and KL regularizers are negligible. There are efficient algorithms for computing sparsemax ($\rho=2$) and $1.5$-entmax; otherwise computing softmax is faster. We use reparametrization tricks in all the VAE experiments: the classical Gaussian one [9], and a $(2-\rho)$-Gaussian reparametrization [39]. We keep computational resources as close as possible between the compared models.
4.1 FYEM for GMM
To instantiate the FYEM algorithm described above to estimate a Gaussian mixture model (GMM), we set $\mathcal{L}_{\theta}(x_{i},z_{i})=-\log\mathcal{N}(x_{i};\mu_{z_{i}},\Sigma_{z_{i}})$ as Gaussian negative log-likelihoods. Setting $\Omega=\Omega_{1}$ recovers the standard EM algorithm for GMM estimation. We also consider other regularizers, $\Omega_{\rho}$ with $p>1$, which yield a new adaptively sparse EM algorithm for GMMs. In Appendix B we review of the baseline EM and hard EM algorithms compared in Figure 1, give details of how the clusters are generated and evaluated in Table 1, and an analysis of the sparsity of E-steps. Figure 2 illustrates how average sparsity and the unuspervised clustering evaluation metrics we use vary with $\rho$.
Figure 1: Comparison of the standard, sparse, and hard versions of EM for GMM-based clustering. Data points are colored by their true labels, with x markers indicating outliers. Black ellipses are level curves of the estimated GMM components, while colored ellipses denote the ground-truth.
Setting $\Omega=\Omega_{\rho}$, the Tsallis negentropy (10) with $\rho>1$, yields the new sparse version of (15):
${\hat{q}_{i}^{\mbox{\scriptsize new}}}=\rho\mbox{-entmax}\Bigl{(}\frac{\pi_{1}^{\rho-1}}{\rho-1}+\log\bigl{(}\mathcal{N}(x_{i};\hat{\mu}_{1},\hat{\Sigma}_{1})\bigr{)},...,\frac{\pi_{K}^{\rho-1}}{\rho-1}+\log\bigl{(}\mathcal{N}(x_{i};\hat{\mu}_{K},\hat{\Sigma}_{K}\bigr{)}\Bigr{)}.$
For $\rho>1$, the output of $\rho\mbox{-entamax}$ can have both zero and non-zero components. When $\rho=1$, Tsallis $\rho$-negentropy becomes the Shannon nentropy and we recover the standard EM algorithm, where all the components of $\hat{q}_{i}^{\mbox{\scriptsize new}}$ are strictly positive. The novel feature of sparse EM is the possible presence of zeros in $\hat{q}_{i}$, meaning that sample $x_{i}$ will not contribute to the update of the parameter estimates of the corresponding components. In standard EM all samples affect the parameter estimates of all the components. In hard EM, each sample only affects the parameters of the “closest" component, as in the case of $K$-means clustering (with the important difference that $K$-means clustering does not estimate component covariances).
Figure 2: We vary $\rho\in\{0.1,0.5,0.9,1.0,1.1,1.5,2.0,3.0\}$ and plot the resulting adjusted mutual information, adjusted Rand index, and silhouette score. If $\rho$ is close to 1, the value of metric is close to its value with classical EM. We quantify the E-step sparsity by counting the number of clusters assigned 0 mixing proportion at the last iteration of each EM run and averaging across samples. Since the total number of clusters is four, the greatest number of components assigned 0 proportion is three. The shaded regions represent standard deviations from the mean values across 5 seeds.
Table 1: Clustering quality metrics (higher is better) for GMM-based clustering using the standard, hard, and sparse (ours) versions of EM; $\rho=2$ in sparse EM. Values are averaged over 5 random seeds. Priors are learned according to Eq. (16).
E-Step Adjusted MI Adjusted RI Silhouette Sc.
Standard EM .606 $\pm$ .012 .531 $\pm$ .005 .345 $\pm$ .032
Hard EM .537 $\pm$ .088 .348 $\pm$ .043 .207 $\pm$ .128
Sparse EM .636 $\pm$ .017 .476 $\pm$ .037 .393 $\pm$ .041
4.2 Fenchel-Young $\beta$-VAE
In Kingma and Welling [33]’s classic VAE a Gaussian variational posterior and a $\mathcal{N}(0,I)$ prior are combined with a product of independent Bernoulli’s as observation model, representing the pixels of MNIST digit images. The standard VI baseline $\beta$-VAE objective is included in in Appendix C. We experiment with truncated parabola $(2-\rho)$-Gaussian [^5] Martins et al. [39] call this family $\beta$-Gaussian. finite-support variational posteriors, comparing the baseline product of Bernoulli’s decoder with a product of binary $\Psi_{2}$ deformed exponential family distributions. The FYELBO objective is
$\sum\limits_{i=1}^{N}\underset{Z\sim q_{i}}{\mathbb{E}}\left[L_{\Psi_{\rho^{\prime}}}(\theta(Z);x_{i})\right]+\beta L_{\Omega_{\rho}}(\eta_{\mathcal{N}(0,\textnormal{I})};q_{i}),$
[…]
Table 2: VAE $\ell_{1}$ reconstruction error on MNIST and Fashion MNIST datasets.
Observation Latent Variable Mnist FashionMnist
Product of independent Bernoulli’s, $\rho^{\prime}=1$ Gaussian, $\rho=1$ 13.272 34.192
Biweight, $\rho=3/2$ 12.061 32.690
Epanechnikov, $\rho=2$ 12.111 32.482
Product of $\Psi_{2}$ sparse binary distributions, $\rho^{\prime}=2$ Gaussian, $\rho=1$ 12.610 32.986
Biweight, $\rho=3/2$ 10.326 25.985
Epanechnikov, $\rho=2$ 9.183 26.186
In Table2 we report the results of computationally efficient combinations of $\Psi$ and $\Omega$ (i.e. strictly comparable with standard VI VAEs) in VAEs applied to MNIST and FashionMNIST. It is noteworthy that sparsity in the encoder and in the encoder separately contribute to improved performance. We report average $l_{1}$ reconstruction error as a means of comparing the different models, since the FY ELBO values themselves are not comparable. In all cases we tune $\beta$ by cross-validation to prevent the posterior from collapsing to the prior.
In Appendix D we report improved $l_{1}$ reconstruction error in a small document model VAE, using sparsity only in the observation model. This can be seen as a generalized Bayesian treatment of sparse observation models, falling under the cases of generalized VI described by Bissiri et al. [7], Khan and Rue [32], Knoblauch et al. [34]. In preliminary experiments with sparse posteriors we did not observe improvements over the baseline NVDM of Miao et al. [40]. This is intuitively reasonable: While the empirical distribution over words is sparse, one latent Multivariate Gaussian associated with each Multinomial (or $\Omega_{\rho}$-sparse) observation through amortized variational inference does not lead to improved reconstruction if its domain is truncated. In the image VAE data, similar patterns of pixels are 0 across images, e.g. on the boarders, while in the document model VAEs there is no word of the vocabulary that ought to be assigned no probability mass in all “topic dimensions”.
5 Conclusion
We introduce Fenchel-Young variational inference (FYVI), expanding the class of variational posteriors to the $(2-\rho)$-deformed exponential family. FYVI exploits a general class of entropic regularizers to obtain sparse discrete and continuous posterior distributions, and distinctly extends classical Bayesian inference compared to previous generalized Bayesian and variational inference proposals like those of Bissiri et al. [7], Knoblauch et al. [34], and Daudel et al. [18]. We focus on cases where the entropic index $\rho$ is greater than $1$, which leads to sparse posteriors previously unexplored in the related literature. Note the nature of sparsity of $(2-\rho)$-exponential family posteriors is qualitatively distinct from the sparse MAP estimates that are obtained through sparsity-inducing priors (§3). We define and lower bound the FY variational free energy (FYVFE), and describe EM-style and amortized VI algorithms for estimating arbitrary generative FYVI models. We derive and experiment with a sparse EM algorithm for Gaussian mixtures, demonstrating better clustering quality than classical and hard versions of EM on synthetic data with overlapping clusters and outliers. In our image FYVAE experiments with continuous finite support posteriors, we obtain the best performance when combining a latent $(2-\rho)$-Gaussian posterior with a sparse support observation model. We also show how to treat sparse observation models from a generalized Bayes perspective.
FYVI can be directly extended to structured latent variables by using SparseMAP [46] posteriors, intuitively sparse analogues to conditional random fields. In future work we plan to extend our analysis and experiments to the non-trivial case of heavy tailed posteriors ($\rho$ < 1). Further, FYVI is a necessary step toward characterizing conjugate FY families and generalized canonical gradient descent, to account for the geometry of the space of deformed exponential family distributions, and to extend information-theoretic generalization bounds that have been studied in the case of SGD [45, 14] to locally reparametrization-robust optimization methods. Another interesting direction is to connect FYVI with variational Rényi objectives [17, 18]. The recently popular $\lambda$-deformation [60, 57, 26], which aims to treat Rényi and Tsallis entropies in a unified way, may provide an illuminating bridge among alternative generalized VI perspectives. Finally, connections with PAC-Bayes analyses of VI like Alquier et al. [3] are also an interesting prospect.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A $\rho$-MaxEnt and relation with Amari’s $\alpha$- and $q$-divergences
$\rho$-deformation definitions.
Defining the Tsallis entropy relies on the definitions of the $\rho$-deformed logarithmic and exponential functions for an entropic index $\rho\in\mathbb{R}\setminus\{1\}$ — this index is called $q$ by Tsallis [53], Amari and Ohara [4], Naudts [43], and $\alpha$ by Martins et al. [39]:
$\log_{\rho}(x):=\frac{x^{1-\rho}-1}{1-\rho},\quad x>0,\quad\exp_{\rho}(x):=[1+(1-\rho)x]_{+}^{\frac{1}{1-\rho}},\quad x\in\mathbb{R}$ (18)
where $\cdot]_{+}:=\max\{\cdot,0\}$, which in the limit $\rho\rightarrow 1$ converge to $\log(x)$ and $\exp(x)$. Note the following relations:
$\log_{\rho}\left(\frac{x}{y}\right)=\frac{\log_{\rho}x-\log_{\rho}y}{1+(1-\rho)\log_{\rho}y},\quad\log_{\rho}(xy)=\log_{\rho}x+\log_{\rho}y-(1-\rho)\log_{\rho}x\log_{\rho}y.$ (19)
$L_{\Omega}$ and related divergences.
Eq. (10) above defines the Tsallis $\rho$-negentropy scaled by $\frac{1}{\rho}$ to preserve consistency with Martins et al. [39] $\Omega_{\rho}(q)=-\frac{1}{\rho}S_{\rho}(q)$ where $S_{\rho}(q)=\frac{1}{1-\rho}\mathbb{E}_{q(z)}\left[q(z)^{\rho-1}-1\right]$ is the Tsallis entropy. Additionally, noting that $\log_{\rho}(\frac{1}{q})+\log_{2-\rho}q=0$
$S_{\rho}(q)=\frac{1}{1-\rho}\mathbb{E}_{q(z)}[q(z)^{\rho-1}-1]=\mathbb{E}_{q(z)}\log_{\rho}\frac{1}{q(z)}=-\mathbb{E}_{q(z)}\log_{2-\rho}q(z).$ (20)
The Tsallis $\rho$-relative entropy, given by Amari and Ohara [4, Eq. 39] is a scalar multiple of Amari and Ohara [4]’s $\alpha$-divergence (equivalent for $\alpha=1-2\rho$):
${D}_{\rho}^{\textnormal{Tsallis}}(q,p)$ $=-\underset{Z\sim q(z)}{\mathbb{E}}\log_{\rho}\frac{p(z)}{q(z)}=-\int q(z)^{\rho}[\log_{\rho}p(z)-\log_{\rho}q(z)]dz$ (21)
$=\frac{1}{1-\rho}[1-\int q(z)^{\rho}p(z)^{1-\rho}dz].$ (22)
Amari and Ohara [4, Eq. 40] propose a $\rho$-divergence, which is a Bregman divergence based on the convex $\rho$-cumulant function, $A_{\rho}$ between two $\rho$-exponential densities with form $\log_{\rho}q=\eta_{q}z-A_{\rho}(\eta_{q})$. Differentiating the $\rho$-cumulant $A_{\rho}$ yields a dual parametrization of the $\rho$-exponential family that is equal to the escort mean $q=\mathbb{E}_{\tilde{q}}z$, where $\tilde{q}(z;\eta)=\frac{q(z;\eta)^{\rho}}{\int q(z;\eta)^{\rho}dz}$ is the escort distribution and the escort expectation parameter $q$ is denoted by the name of the distribution. Amari and Ohara [4] show that the $\rho$-divergence is a “conformal transformation” of ${D}_{\rho}^{\textnormal{Tsallis}}$. Note also that if we multiply Amari’s $\alpha$-divergence by $\frac{\rho}{\int p_{i}^{\rho}}$ (for $\rho=\frac{1-\alpha}{2}$) we retrieve the $\rho$-divergence:
$\begin{split}D_{\rho}(q,p)&=\frac{\rho}{\int q_{i}^{\rho}}\underbrace{\frac{1}{\rho}D^{\textnormal{Tsallis}}_{\rho}(q,p)}_{\textnormal{Amari's }\alpha\textnormal{-divergence}}=\frac{-\int q_{i}(\log_{\rho}p_{i}-\log_{\rho}q_{i})}{\int q_{i}^{\rho}}=\underset{{\color[rgb]{1,0,0}\tilde{q}}}{\mathbb{E}}[\log_{\rho}q_{i}-\log_{\rho}p_{i}]\\ &=\underset{{\color[rgb]{1,0,0}\tilde{q}}}{\mathbb{E}}[\eta_{q}z-A_{\rho}(\eta_{q})-(\eta_{p}z-A_{\rho}(\eta_{p}))]=\eta_{q}q-A_{\rho}(\eta_{q})-(\eta_{p}q-A_{\rho}(\eta_{p}))\\ &=A_{\rho}(\eta_{p})-A_{\rho}(\eta_{q})-\underbrace{q}_{=\nabla A_{\rho}(\eta_{q})}(\eta_{p}-\eta_{q})=\textnormal{Bregman}_{\color[rgb]{1,0,0}{A_{\rho}}}(\eta_{p},\eta_{q})\\ &=A^{*}_{\rho}(q)+A_{\rho}(\eta_{p})-\eta_{p}q=L_{\color[rgb]{1,0,0}A^{*}_{\rho}}(\eta_{p};q).\end{split}$ (23)
For the expressions in the last line we use the Legendre duality $\eta_{q}q-A_{\rho}(\eta_{q})=A^{*}_{\rho}(q)$. $L_{\Omega}$ coincide with the Bregman divergence generated by a strictly convex and differentiable $\Omega$, when in addition $\Omega$ is a smooth Legendre type function, meaning that its gradient explodes in the boundary of its domain [10, §3.2].
The term $L_{A^{*}_{\rho}}(\eta_{p};q)$ is a divergence between distributions $p$ and $\tilde{q}$, which is different from the FY loss/divergence that we use in the main text between prior canonical parameters $\eta_{p}$ and variational mean parameters $q$ since the dual of the $\rho$-cumulant $A^{*}_{\rho}$ is different from the $\Omega$ that we use in the FY term. We discuss the difference below.
Maximum Tsallis $\rho$-Entropy.
There are two relevant approaches to derive the same $\rho$-exponential family: maximize the $(2-\rho)$-entropy subject to linear expectation constraints, or maximize the $\rho$-entropy subject to escort expectation constraints. Amari and Ohara [4] follow the latter approach and are interested in cases where the two densities $q,p$ belong to the $\rho$-expfam that maximizes the Tsallis $\rho$-entropy subject to escort distribution constraints. In this paper, we focus on the $(2-\rho)$-exponential family of distributions that arises from the Tsallis $\rho$-MaxEnt principle subject to standard linear expectation constraints, following Naudts [43] and Martins et al. [39]. These densities are the $\Omega_{\rho}$-regularized prediction maps $p(z;\eta)=\exp_{2-\rho}[\eta z-A_{\rho}(\eta)]$ (c.f. Martins et al. [39, Def. 2]). Overloading $\eta$ to denote both a function and canonical parameters, the scoring function is $\eta(z)=\eta z$. The FY loss in Eq. (6) above, induced by the negentropy $\Omega$ essentially compares two members of the $(2-\rho)$-exponential family: $L_{\Omega}(\eta;q)=\Omega_{\rho}(q)-\Omega_{\rho}(p(z;\eta))-\eta^{\textnormal{T}}(\mathbb{E}_{p_{\eta}}[z]-\mathbb{E}_{q}[z])$ (c.f. Martins et al. [39, Eq. 14]).
Starting with regularized prediction maps like Martins et al. [39, Def. 2] the index $\rho$ ($\alpha$ in their notation) of the Tsallis $\rho$-entropy is specified first, and then the associated deformed family is defined using $\exp_{2-\rho}$, based on the $\rho$-MaxEnt problem under standard constraints [39, Eq. 10]. On the other hand, if the index $\rho$ is used to write a $\rho$-exponential family first, then the associated Tsallis entropy that is maximized, again with standard constraints, is the $(2-\rho)$-entropy (see also Naudts [44, Chapters 7, 8] and Wada and Scarfone [55], Wada et al. [56].) Amari and Ohara [4] use the index $\rho$ to write the $\rho$-deformed family using $\exp_{\rho}$. They concentrate on the “integration constant” that they call the $\rho$-log partition function (their $\psi$ is the same as the $A_{\alpha}$ of Martins et al. [39], given correct mapping of indices; we use $A_{\rho}$ above). Then, Amari and Ohara [4] derive the dual of $\psi$ which they call $\rho$-negentropy. Their $\rho$-divergence is the Bregman divergence based on the $\rho$-log partition, which is a useful quantity. Martins et al. [39, Prop. 10] also defines useful quantities: they derive the dual of the Tsallis negentropy to use in the FY loss. Suppose that we use $\rho$ as Amari and Ohara [4] for the $\rho$-exponential. Then Martins et al. [39, Prop. 9] describes how to take the associated $(2-\rho)$-Tsallis negentropy as $\Omega$ and derive its dual $\Omega^{*}$, which is not the log normalizer $A_{\rho}$.
Proposition 1 (Fenchel-Young VI and Max Tsallis Entropy Equivalence).
Let $\mathcal{Q}$ be a convex set of probability distributions over a space $\mathcal{Z}$, $\ell(x;z)$ a measurable loss function, and $\eta:\mathcal{Z}\rightarrow\mathbb{R}$ a scoring function (possibly a log-prior). Let $\Omega:\mathcal{Q}\to\mathbb{R}\cup\{+\infty\}$ be a proper, lower semi-continuous, convex regularizer that induces a Fenchel-Young loss:
$L_{\Omega}(\eta;q):=\Omega(q)+\Omega^{*}(\eta)-\mathbb{E}_{Z\sim q}[\eta(Z)].$
Then the variational problem
$q_{x}^{\star}=\arg\min_{q\in\mathcal{Q}}\mathbb{E}_{Z\sim q}[\ell(x;Z)]+L_{\Omega}(\eta;q)$ (24)
is equivalent to a generalized Tsallis maximum entropy problem under standard expectation constraints, where the regularizer $\Omega$ is chosen as the negative Tsallis $\rho$-entropy for $\rho>1$:
$\Omega_{\rho}(q)=-\frac{1}{\rho}S_{\rho}(q)=\frac{1}{\rho}\frac{1}{\rho-1}\left(1-\int q(z)^{\rho}dz\right).$
Moreover, the solution $q_{x}^{\star}$ belongs to the $(2-\rho)$-exponential family, i.e.,
$q_{x}^{\star}(z)$ $=\left[1+(1-\rho)\left(\eta(z)-\ell(x;z)-A_{\rho}\right)\right]_{+}^{\frac{1}{\rho-1}}$ (25)
$=\exp_{2-\rho}\left[1+\eta(z)-\ell(x;z)-A_{\rho}\right],$ (26)
where $A_{\rho}$ is a normalization function.
Proof.
The variational problem
$q_{x}^{\star}=\arg\min_{q\in\mathcal{Q}}\mathbb{E}_{Z\sim q}\left[\ell(x;z)\right]+L_{\Omega}(\eta;q)$
defines the $\Omega$-regularized prediction map with scoring function $\eta(z)-\ell(x;z)$
$q_{x}^{\star}(z)=\Pi_{\Omega}(\eta(z)-\ell(x;z))=\arg\max_{q\in\mathcal{Q}}\mathbb{E}_{Z\sim q}\left[\eta(z)-\ell(x;z)\right]-\Omega(q)$
that by definition maximizes the $\rho$-Tsallis entropy subject to an expectation constraint. Using Martins et al. [39, Prop. 8] we can express $q_{x}^{\star}(z)$ as a member of the $(2-\rho)$-exponential family. ∎
Appendix B Clustering with FY EM
Standard EM.
The FY EM algorithm described in $\S$3.3 recovers standard EM for GMMs when:
  • $x_{1},...,x_{N}\in\mathbb{R}^{d}$ are $d$-variate observations and the latent variables $z_{1},...z_{N}\in\mathcal{Z}=\{1,...,K\}$ indicate to which component each observation $x_{i}$ belongs;
  • $q_{i}=\bigl{(}q_{i}(1),...,q_{i}(K)\bigr{)}\in\!\Delta(\mathcal{Z})$ is the variational approximation of the distribution of $z_{i}$;
  • $\eta\in\mathbb{R}^{K}$ contains the logarithms of the prior probabilities of the $K$ components, denoted $\pi_{1},...,\pi_{K}$, i.e., $\eta_{1}=\log\pi_{1},...,\eta_{K}=\log\pi_{K}$;
  • $\ell(x_{i},z,\theta)=-\log\mathcal{N}(x_{i};\mu_{z},\Sigma_{z})$, for z=1,,Kz=1,...,K, where $\mu_{z}$ and $\Sigma_{z}$ are, respectively, the mean and covariance of component $z$; and
  • $\Omega(q)=-H(q)$, thus $\nabla\Omega^{*}=$ softmax.
The E-step in (15) then takes the familiar form
$\hat{q}_{i}^{\mbox{\scriptsize new}}(z)=\mbox{softmax}\bigl{(}\eta_{z}+\log\mathcal{N}(x_{i};\hat{\mu}_{z},\hat{\Sigma}_{z})\bigr{)}=\pi_{z}\;\mathcal{N}(x_{i};\hat{\mu}_{z},\hat{\Sigma}_{z})\biggl{(}\sum_{j=1}^{K}\pi_{j}\;\mathcal{N}(x_{i};\hat{\mu}_{j},\hat{\Sigma}_{j})\biggr{)}^{-1}\!,$
for i=1,,N,i=1,...,N, z=1,,Kz=1,...,K, where $\hat{\mu}_{z}$, $\hat{\Sigma}_{z}$ are the current parameter estimates. This is the E-step of standard EM for GMM estimation [23]. The M-step corresponds to maximum weighted log-likelihood estimation, yielding, for $z_{i}=1,...,K$,
$\hat{\mu}_{z}^{\mbox{\scriptsize new}}$ $=\frac{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}(z)\;x_{i}}{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}(z)},\quad\hat{\Sigma}_{z}^{\mbox{\scriptsize new}}=\frac{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}(z)\;(x_{i}-\hat{\mu}_{z}^{\mbox{\scriptsize new}})(x_{i}-\hat{\mu}_{z}^{\mbox{\scriptsize new}})^{\top}}{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}(z)},$
$\hat{\eta}_{z}^{\mbox{\scriptsize new}}$ $=\log\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}(z)-\log N.$
Hard EM.
Setting $\Omega(q)=0$, i.e., removing the regularizer, yields a different E-step:
$\hat{q}_{i}^{\mbox{\scriptsize new}}\!=\mbox{argmax}\bigl{(}\pi_{1}\;\mathcal{N}(x_{i};\hat{\mu}_{1},\hat{\Sigma}_{1}),...,\pi_{K}\;\mathcal{N}(x_{i};\hat{\mu}_{K},\hat{\Sigma}_{K})\bigr{)},$
where $\mbox{argmax}(u_{1},...,u_{K})$ of a tuple of $K$ numbers returns a vector with $1/m$ in the entries with the same indices as the $m$ maxima and zero everywhere else. If the maximum is unique, this is just the one-hot representation of the maximizer. This hard version of EM coincides with the classification EM (CEM) algorithm proposed by Celeux and Govaert [13]. The M-step is the same as in standard EM.
Experimental Details.
We generate $1000$ samples uniformly from four bivariate Gaussians with means at [1,1][-1,-1], $[0,0]$, $[1,1]$, and [1,1][1,-1]. The components overlap significantly, with covariance matrices 0.11I0.11\,I, 0.5I0.5\,I, 0.7I0.7\,I, and 0.9I0.9\,I. Additionally, we sample $100$ uniformly distributed random samples in $3,3]^{2}$ to simulate the presence of outliers, yielding a total of $1100$ data points. We compare the three EM variants described —standard EM, hard EM, and sparse EM— on unsupervised clustering of this dataset. The number of components in each model is set to $4$ and the initial parameters (means and covariances) are randomly initialized from a uniform distribution in [0, 0.1][0,\,0.1]. Each algorithm runs $200$ iterations, which we verified is more than enough to satisfy a tight convergence criterion.
We evaluate the quality of unsupervised clustering of each algorithm using three metrics: adjusted mutual information [54], adjusted Rand index [48], and silhouette score [51]. Each model is trained and evaluated across five different random seeds, and the average and standard deviation of the metrics are reported in Table 1, showing that sparse EM outperforms standard and hard EM in two of three standard metrics. See also Figure 1 for a qualitative illustration. In Figure 2 we vary $\rho=0.1,0.5,0.9,1.0,1.1,1.5,2.0,3.0$ and evaluate the same three metrics when clustering with the corresponding EM. In the right most image of Figure 2 we verify that the sparsity of the E-step corresponds to the value of $\rho$ for the same values of $\rho$. We quantify the E-step sparsity by counting the number of clusters assigned 0 mixing proportion at the last iteration of each EM run and averaging across samples. Since the total number of clusters is four, the greatest number of components assigned 0 proportion is three.
Appendix C MNIST and FashionMNIST
The objective maximized by standard amortized VI w.r.t. $\theta$ (decoder parameters) and $\phi$ (encoder parameters), for a set of $N$ images is
$\sum\limits_{i=1}^{N}\underset{Z\sim q_{i}}{\mathbb{E}}\bigl{[}\log p_{\theta}(x_{i}|Z)\bigr{]}-\beta D_{\text{\small KL}}(q_{i}\|p_{Z}),$ (27)
where $q_{i}=\mathcal{N}\bigl{(}\mu_{\phi}(x_{i}),\mbox{diag}\bigl{(}\sigma^{2}_{\phi}(x_{i})\bigr{)}\bigr{)}$ and
$p_{\theta}(x_{i}|z_{i})\sim\textnormal{Bernoulli}(\sigma(f_{\theta}(z_{i}));$
$\sigma$ denotes the sigmoid function, and $f_{\theta}$ a neurally parametrized function.
Experimental Details.
[…]
We report the $\ell_{1}$ reconstruction error in Table 2, i.e. the absolute distance between the model’s reconstructed samples and the original input images, evaluated on the test split for both MNIST and FashionMNIST datasets. [^6] MNIST is Public Domain, and FashionMNIST is under an MIT License [58]. The reconstruction error is the only standard and interpretable metric that is comparable across the VAE alternatives we consider. The FY ELBO has a log scale only in the standard VAE case, which is lost when either the observation or latent variable model is based on a Tsallis entropy with $\rho>1$. The first row of Table 2 corresponds to the standard VAE with a Gaussian latent variable. The Biweight and Epanechnikov densities are limited support distributions of the latent variable, with Tsallis entropies corresponding to $\rho=3/2$ $\rho=2$ respectively. The best reconstruction errors are obtained with latent $(2-\rho)$-Gaussian encoders combined with 2-entmax observation models.
Appendix D Document Fenchel-Young $\beta$-VAE
Similarly with one of the cases we compared in the image VAE, namely the sparsemax observation model combined with a Gaussian latent variable in the previous section, we experiment with a FY $\beta$-VAE applied to unsupervised word embeddings to directly model the sparsity of the empirical distribution. In this experiment we focus on tuning the level of sparsity in the observation model rather than using only $\rho^{\prime}=2$. The baseline neural variational document model (NVDM) proposed by Miao et al. [40] is a standard VAE (with prior $p_{Z}=\mathcal{N}(0,I)$): it models the joint probability of observed word counts (or proportions) in a document $x_{i}$, i=1,,Ni=1,...,N, and of continuous latent semantic embeddings $z_{i}$, by optimizing the same objective as (27) and replacing the multivariate Bernoulli with a multinomial observation model,
$p_{\theta}(x_{i}|z_{i})\sim\textnormal{Multinomial}(\textnormal{softmax}(Wz_{i})),$
where $\theta=W\in\mathbb{R}^{|D|\times K}$, $K$ is the dimension of the latent multivariate Gaussian, and $|D|$ is the size of the observed vocabulary. The model parameters are learned by stochastic gradient backpropagation.
Our FY $\beta$-VAE model uses the same multivariate Gaussian latent $q$ and prior $p_{Z}$, combined with a FY observation model $L_{\Psi}(\theta(z);x)$:
$J(\phi,\theta)=\sum\limits_{i=1}^{N}\underset{Z\sim q}{\mathbb{E}}\left[L_{\Psi}(\theta(Z);x_{i})\right]+\beta D_{\text{\footnotesize KL}}(q_{i}\|p_{Z}),$
where $\Psi$ is a Tsallis $\rho^{\prime}$-negentropy, dom($\Psi)=\Delta_{|D|-1}$, and $\rho^{\prime}$-entmax replaces softmax in the $L_{\Psi}$ decoder. Unlike the standard VAE formulation, the deformed exponential family distribution parametrized by $\theta(z)$ can assign zero probability to vocabulary categories that do not occur in a given document, directly modeling the sparse empirical distribution.
Experimental Details. We use the 20NewsGroups [^7] http://qwone.com/~jason/20Newsgroups/, [35]. dataset, which consists of 11,314 training samples and 7,531 test samples (documents, represented as bags of words). We assume a decoder parameterizing a sparse discrete distribution over the vocabulary generates all the words in a document independently, given a continuous Gaussian latent variable $Z$, i.e. a dense, continuous document representation. Following Miao et al. [40], we assume that each dimension in the latent space represents a “topic”, which is a simplification of the classical discrete latent variable topic model of [8].
The first 1,000 test samples were used as the validation set for hyperparameter tuning. We perform a search over entmax $\rho^{\prime}$, which controls the decoder’s sparsity, learning rate, batch size, number of epochs, and the number of topics using Optuna [1] across 100 trials for both the baseline and FY NVDM models. Finally, we rerun the experiment with the best hyperparameters using five different random seeds and report the reconstruction error (RE) defined by the $\ell_{1}$ distance between the empirical data distribution and the learned observation model distribution. For the baseline NVDM model over 100 Optuna trials the hyperparameters that yield the best reconstruction error of 1.484 reported in Table 3 are batch size = 64, number of alternating epochs updating the decoder and encoder parameters = 1. For the Fenchel-Young $\beta$-VAE over 100 Optuna trials the hyperparameters that yield the best reconstruction error of 1.384 reported in Table 3 are $\rho^{\prime}=1.314$, batch size = 1024, number of alternating epochs updating the decoder and encoder parameters = 3. In both cases $\beta=0.01$, learning rate $5\times 10^{-5}$, number of training epochs = 500, and dimension of the latent Gaussian = 50. Similarly to the image VAE experiments we use $\beta=0.01$ ensures that the regulazation term is non trivially small. Both models converge based on training objective curves, and in preliminary experiments we empirically verified that for $\rho^{\prime}$ close to 1, training and validation behavior was similar to the NVDM baseline. Again the objectives are are not numerically comparable to each other, nor is expected $\Psi_{r}^{\prime}$ term with the log scale regularization term. Again $\ell_{1}$ reconstruction error is a metric that is comparable between the NVDM and the FY NVDM.
Our sparse support observation model FYVAE outperforms the NVDM baseline method [40] by 10.68% in absolute reconstruction error, as reported in Table 3.
Table 3: Average reconstruction error over 5 seeds for each model normalized by the number of examples in the test set (6505).
Model Average RE
NVDM 1.4984 ± 0.0009
FY VAE 1.3384 ± 0.0008
Figure 3: Topic coherence scores for the standard NVDM and the FY VAE models.
Figure 3 shows average coherence scores, measured by the pairwise similarity between embedding representations of the top-$k$ words for each latent “topic” dimension, using a sentence-transformer model [49]. [^8] We use the all-mpnet-base-v2 model, which is available at huggingface.co/sentence-transformers/all-mpnet-base-v2, under Apache-2.0. FYVAE consistently achieves better or similar coherence scores across varying values of $k$ (number of words) than the NVDM model, confirming that the latent document representations learned by FYVAE are more semantically coherent and better aligned with meaningful word groupings.