### v3

# Fenchel-Young Variational Inference

Abstract
Bayesian inference has a long history in machine learning but struggles in modern applications where exact inference is both infeasible and fragile due to the absence of well-specified priors and likelihoods. Variational inference with generalized objectives and approximating families aims to address these issues, but it is not always easy to adapt existing algorithms to the generalized setting. We present a new framework called Fenchel-Young variational inference (FYVI), which couples tightly the variational family to the objective. This coupling leads to easy and natural generalizations of existing variational methods to general posterior forms, which we exemplify through the expectation-maximization algorithm for latent variable models. We focus on a class of objectives based on Tsallis entropies which allow us to obtain generalized posteriors with support smaller than that of the prior. In several controlled experiments where modeling with sparse posteriors is intuitively appealing, we show that our generalized algorithms outperform established counterparts.
1 INTRODUCTION
Bayesian inference has long played a central role in machine learning, for example, as put forward in the seminal textbooks of bishop2006pattern and murphy2012machine. Despite its foundational role, the most modern applications in deep learning struggle to make use of Bayesian inference due to lack of well-specified priors, likelihoods, as well as the sheer scale of the problems making exact inference intractable. Pursuing early ideas of Zellner1988, recent work by knoblauch2022optimization and JMLR:v24:22-0291, as well as PAC-Bayesian approaches (see the recent survey of alquier2024user), advocate for an optimization-centric variational view of Bayesian inference. This view leads to models and algorithms with desirable Bayesian characteristics but which do not necessarily aim to execute or to approximate Bayes rule in a well-specified model.
JMLR:v24:22-0291 unify diverse, well-established Bayesian inference algorithms and modern deep learning methods under a common paradigm: Using natural-gradient descent to find exponential family posteriors over the unknown parameters or latent variables that strike a balance between a low expected risk and a small Kullback-Leibler divergence (KLD) from a given prior. This line of work led to recent successes, for instance, in Bayesian deep learning [vogn, ivon], large language model finetuning [cong2025improving], language generation [daheim2025uncertainty] and federated learning [mollenhoff2025federated]. However, it remains an open question how such formulations can be naturally extended from Kullback-Leibler divergences and exponential family posteriors to more general divergences and families which may work better in modern settings.
In this paper, we propose a new framework called Fenchel-Young variational inference (FYVI) which addresses this problem, and provides new generalized notions of FY free energy, FY evidence, FY evidence lower bound, and FY posterior. FYVI generalizes the formulations of [knoblauch2022optimization, JMLR:v24:22-0291], leading to a strictly larger class of statistical families which include generalized posteriors with limited (sparse) support, as well as heavy-tailed posteriors. A key property of the new framework is that it couples tightly the variational family to the variational objective, which we argue is the right modeling that leads to natural generalizations of existing algorithms to general posterior forms. This is where FYVI departs from knoblauch2022optimization who analyze arbitrary combinations of divergences, losses, and exponential variational families, and generalizes the modeling of JMLR:v24:22-0291 who are restricted to KLD and exponential families.
The starting point for our new FYVI framework is the fact that the Bayesian posterior is the solution of an optimization over probability distributions where the objective is the expected log-likelihood plus the Kullback-Leibler divergence of a variational posterior from the prior [donsker1975asymptotic, csiszar1975divergence, Zellner1988]. The KLD formulation is equivalent to a maximum entropy problem that favors posteriors with high Shannon entropy [Jordan1999], giving a privileged position to exponential families [Grunwald_Dawid], following from the well-known fact that these are maximum Shannon entropy distributions under linear expectation constraints. The main idea of FYVI is to consider alternative optimality criteria that exploit the class of Fenchel-Young (FY) expressions [Blondel2020, martins2022sparse] and thereby extend the variational formulations of generalized posterior inference [bissiri2016general, knoblauch2022optimization, JMLR:v24:22-0291]. We remark that while Fenchel-Young losses have been used in the past for empirical risk minimization and discriminative learning [Blondel2020, martins2022sparse], our FYVI framework is radically different, as it uses FY for generalized Bayesian learning.
If we replace the KLD with an FY term in variational objectives, we obtain generalized deformed exponential family posteriors, which unlike exponential families may have variable support. We will focus on the particular case of Tsallis entropy FY objectives, which subsume the KLD. In particular, we consider sparse posteriors, but note that heavy-tailed posteriors are an important alternative case that also falls under the proposed framework. We derive a new EM-style [dempster1977maximum] algorithm called FYEM for parameter estimation, and describe gradient backpropagation for amortized inference with deformed exponential family posteriors. Either of these estimation methods could be used to fit arbitrary generative models specifiable as instances of FYVI, and the proposed algorithms are natural generalizations of well-known KLD-based counterparts.
We experiment with the FYEM algorithm applied to Gaussian mixture model (GMM) [murphy2022probabilistic] estimation, and two FY variational autoencoders (VAEs) [Kingma2014, rezende2014stochastic], one applied to images and the other to documents. The FYEM algorithm for GMMs recovers both classical EM and hard EM as special cases and yields new generalizations, for instance, an adaptively sparse E-step. In the image FYVAE we obtain the best performance when we use a sparse posterior, as also observed in the concurrent work [qin2024kernel]. We also study the effect of changing the posterior form on a neural variational document model (NVDM) [miao2016neural]. All our experiments suggest that our generalized algorithms can outperform established counterparts.
2 BACKGROUND
Definitions and Notation.
[…]
2.1 Variational View of Bayesian Inference
Standard Bayesian Inference.
The Bayesian posterior $p_{Z|X}$ can be expressed in variational form as the distribution in $\Delta(\mathcal{Z})$ which strikes a balance between a small KLD from the prior $p_{Z}$ and a high expected log-likelihood $\log p_{X|Z}$, given some observation $x\in\mathcal{X}$ [Thanh2021, csiszar1975divergence, donsker1975asymptotic, Zellner1988]. Formally,
${p_{Z|X}(\cdot|x)}$
$\quad=\arg\!\!\max_{q\in\Delta(\mathcal{Z})}\,\underset{z\sim q}{\mathbb{E}}\bigl[\log p_{X|Z}(x|z)\bigr]\!-\!D_{\textnormal{KL}}(q\|p_{Z})$ (1)
$\quad=\arg\!\!\max_{q\in\Delta(\mathcal{Z})}\,\underset{z\sim q}{\mathbb{E}}\bigl[\log p_{X,Z}(x,z)\bigr]\!+\!H(q).$ (2)
The optimum value of (2), attained by $q=p_{Z|X}(\cdot|x)$, is equal to the log-evidence:
$\mathbb{E}_{z\sim p_{Z|X}(\cdot|x)}\bigl[\log p_{X,Z}(x,z)\bigr]+H\bigl(p_{Z|X}(\cdot|x)\bigr)$ (3)
$\qquad=\log p_{X}(x).$
More often than not the posterior and evidence cannot be easily obtained, which has stimulated decades of work on approximate Bayesian methods.
Variational Inference.
The goal of variational inference (VI) [Blei2017, Jordan1999] is to approximate the posterior and bound the evidence when these are intractable, by confining $q$ to a tractable family: $\mathcal{Q}\subseteq\Delta(\mathcal{Z})$. The familiar VI problem is
$q_{x}^{\star}=\arg\max_{q\in\mathcal{Q}}\,\underset{z\sim q}{\mathbb{E}}\bigl[\log p_{X,Z}(x,z)\bigr]+H(q),$ (4)
and its optimal value is the evidence lower bound (ELBO): ${\mathbb{E}_{z\sim q_{x}^{\star}}}\bigl[\log p_{X,Z}(x,z)\bigr]+H\bigl(q_{x}^{\star}\bigr)\leq\log p_{X}(x)$. The negative of (4) is called the variational free energy.
Generalized VI.
JMLR:v24:22-0291 extend (4) to loss-based posteriors using bissiri2016general’s generalization of classical VI in order to derive as special cases much of the core catalog of modern machine learning algorithms and to propose new ones. The extension consists of replacing the negative log-likelihood with a loss function $\ell(x;z)$:
$q_{x}^{\star}$ $=\arg\min_{q\in\mathcal{Q}}\underset{z\sim q}{\mathbb{E}}\left[\ell(x;z)\right]+D_{\text{KL}}(q\|p_{Z})$ (5)
$=\arg\min_{q\in\mathcal{Q}}\underset{z\sim q}{\mathbb{E}}\left[\ell(x;z)-\log p_{Z}(z)\right]-H(q),$ (6)
where the notation $q_{x}^{\star}$ highlights the dependence on the observations $x$. JMLR:v24:22-0291 restrict $\mathcal{Q}$ to be a regular minimal exponential family, that is, $\mathcal{Q}=\{q(z)=h(z)\exp\bigl(\langle\lambda,t(z)\rangle-A(\lambda)\bigr),\mbox{for }\lambda\in\Lambda\}$, where $\Lambda$ is the parameter space, $t$ is a canonical sufficient statistic, $h$ is the base measure (independent of $\lambda$), and $A(\lambda)$ is the log-partition function. In a similar vein, knoblauch2022optimization propose generalized VI (GVI), by additionally considering divergences other than KLD, and introducing the rule of three, which refers to the choice of the three elements required for GVI: A loss function, a divergence measuring the deviation of variational posterior from prior, and a set of permissible solutions $\mathcal{Q}$.
2.2 Fenchel-Young Losses
In a previously seemingly unrelated line of research, Blondel2020 propose Fenchel-Young (FY) losses, which generalize the ubiquitous cross-entropy loss, and can be seen as mixed space Bregman divergences (see also [amari2010information, amari2011geometry]). FY losses have been used for supervised problems, but not in any kind of Bayesian contexts like VI or latent variable models. Here we use FY expressions in a fundamentally different way since we do not use them to evaluate discrepancy between the score of some distribution and data. Our notation below treats both discrete and continuous domains in a unified way.
Let $\Omega:\Delta(\mathcal{Z})\rightarrow\bar{\mathbb{R}}$ be a lower semi-continuous (l.s.c.) proper convex function, and let $\Omega^{*}$ be its Legendre-Fenchel conjugate [^1] Given a function $f:\mathcal{U}\rightarrow\bar{\mathbb{R}}$, its Fenchel conjugate $f^{*}:\mathcal{U}^{*}\rightarrow\bar{\mathbb{R}}$ is defined as $f^{*}(x)=\sup_{u\in\mathcal{U}}(\langle x,u\rangle-f(u)),$ where $\mathcal{U}^{*}$ is the dual space of $\mathcal{U}$. If $f$ is convex l.s.c., biconjugation holds: $f^{**}=f$. Assuming differentiability and that the supremum is attained, the maximizing argument in the definition of $f^{*}$ is $\arg\max_{u\in\mathcal{U}}(\langle x,u\rangle-f(u))=\nabla f^{*}(x)$. [Borwein, Bauschke]. And let $\eta:\mathcal{Z}\rightarrow\mathbb{R}$ be a score function that assigns a real-valued score to each element of $\mathcal{Z}$. From a Bayesian perspective $\eta$ can be viewed as a log-prior. For $q\in\Delta(\mathcal{Z})$, over discrete or continuous $\mathcal{Z}$, the FY loss induced by $\Omega$ is a measure of how incompatible $\eta$ and $q$ are:
$D_{\Omega}(\eta;q)$ $=\Omega^{*}(\eta)-\langle q,\eta\rangle+\Omega(q)$ (7)
$=\Omega^{*}(\eta)-\mathbb{E}_{z\sim q}[\eta(z)]+\Omega(q),$ (8)
where the second equality is using the definition of the dual pairing $\langle q,\eta\rangle=\mathbb{E}_{z\sim q}[\eta(z)]$. When $\Omega$ is the Shannon negentropy $D_{\Omega}(\eta;q)=D_{\mathrm{KL}}(q||\exp(\eta)/\int\exp(\eta(z))dq(z))$. An FY loss is always non-negative, due to the Fenchel-Young inequality [Bauschke],
$\Omega^{*}(\eta)$ $=\sup_{q\in\Delta(\mathcal{Z})}\,\langle q,\eta\rangle-\Omega(q)\;$ (9)
$\Rightarrow\;\Omega^{*}(\eta)-\langle q,\eta\rangle+\Omega(q)\geq 0.$ (10)
Furthermore, $D_{\Omega}(\eta;q)$ is zero if and only if $q=\arg\max_{q\in\Delta(\mathcal{Z})}\mathbb{E}_{z\sim q}[\eta(z)]-\Omega(q)$. This defines the $\Omega$-regularized prediction map $\Pi_{\Omega}:\mathcal{F}\rightarrow\Delta(\mathcal{Z})$,
$\Pi_{\Omega}(\eta):=\arg\max_{q\in\Delta(\mathcal{Z})}\mathbb{E}_{z\sim q}[\eta(z)]-\Omega(q),$ (11)
where $\mathcal{F}\subseteq\mathbb{R}^{\mathcal{Z}}$ is the set of functions for which the maximizer exists and is unique [Blondel2020], and we exclude $\Omega$ for which $\mathcal{F}$ is the empty set. A trivial but important property of $\Pi_{\Omega}$ is invariance under global shifts of the argument: $\Pi_{\Omega}[\eta+C]=\Pi_{\Omega}[\eta]$, for any constant $C.$
[…]
We choose $\Omega$ to be the Tsallis negentropy [tsallis2009introduction],
$\Omega_{\rho}(q)=\left\{\begin{array}[]{ll}\hskip-4.26773pt\mathbb{E}_{z\sim q}[q(z)^{\rho-1}-1]/\bigl(\rho^{2}-\rho\bigr),&\text{$\rho\neq 1$,}\\ \hskip-4.26773pt\mathbb{E}_{z\sim q}[\log q(z)],&\text{$\rho=1$,}\end{array}\right$. (14)
where $\rho>0$ is a parameter called the entropic index. Tsallis negentropies are continuous in $\rho$ and coincide with the Shannon negentropy when $\rho=1$. For finite $\mathcal{Z}$ the $\Omega_{\rho}$-regularized prediction map $\Pi_{\Omega_{\rho}}$ is called $\rho$-entmax [peters2019sparse] and recovers softmax and sparsemax, for $\rho=1$ and $\rho=2$ respectively. For any $\rho>1$, the $\Omega_{\rho}$-regularized prediction map will be sparse, with higher values of $\rho$ giving distributions with increasingly sparse supports.
In Appendix A we describe the relationship between Tsallis negentropy FY losses and the $\alpha$- and $q$-divergences of amari2011geometry. These $\alpha$-divergences are connected to Tsallis statistics, and are related to but distinct from the divergences associated with Rényi entropies, which, somewhat confusingly, are also often called $\alpha$-divergences; we do not address Variational Rényi approaches in this paper [li2016renyi, minka2005divergence, knoblauch2022optimization, daudel2023alpha, daudel2023monotonic, guilmeau2023variational].
3 FENCHEL-YOUNG VARIATIONAL INFERENCE
3.1 Fenchel-Young Variational Posteriors
FYVI problems are built by substituting the Kullback-Leibler divergence in (5) with the FY discrepancy induced by a regularizer $\Omega$ with respect to a log-prior $\eta=\log p_{Z}$. This yields a new class of generalized posterior distributions of the following form:
$q_{x}^{\star}$ $=\arg\min_{q\in\mathcal{Q}}\,\underset{z\sim q}{\mathbb{E}}\left[\ell(x;z)\right]+D_{\Omega}(\eta;q)$ (15)
$=\arg\min_{q\in\mathcal{Q}}\,\underset{z\sim q}{\mathbb{E}}\left[\ell(x;z)-\eta(z)\right]+\Omega(q).$ (16)
We first consider the case where $\mathcal{Q}=\Delta(\mathcal{Z})$ is the space of all probability distributions.
We focus on FYVI with Tsallis $\rho$-negentropies $\Omega_{\rho}$ as regularizers, which for $\rho>1$ naturally induce variational posteriors with adaptively sparse support, that is, such that the support may be strictly smaller than $\mathcal{Z}$ and depends on observations $x$. In contrast, when $\rho<1$ the variational posteriors would be denser (closer to uniform) than those obtained with standard Shannon entropy regularization ($\rho=1$); we do not consider this case, and leave it for future work. Figure 1 illustrates the appeal of sparse posteriors on a 1D logistic regression problem: Due to the compact support, the approximate posteriors with $\rho=2$ and $\rho=1.5$ are closer to the MAP solution while still capturing uncertainty around it. Our FYVI framework gives us a new tool to control how close we want to be to the MAP solution, which we expect to be helpful in deep learning.
By definition, exponential families ($\rho=1$) have support that cannot depend on their parameters [Barndorff-Nielsen]. Consequently, due to their variable support, generalized posteriors obtained through (16) where $\Omega$ is a Tsallis $\rho$-negentropy with $\rho>1$, cannot belong to an exponential family. This fact implies that contrary to knoblauch2022optimization’s rule of three suggestion the variational family $\mathcal{Q}$ and the divergence should not naively be chosen independently of each other. For example, adopting a divergence that may yield generalized posteriors with adaptive support is incompatible with a variational family in any exponential family. In Appendix A, following similar derivations as Prop. 8 in martins2022sparse, we show that for $\Omega=\Omega_{\rho}$ the optimization in (16) is equivalent to a maximum Tsallis $\rho$-entropy problem. The solution to that problem belongs to a $(2-\rho)$-exponential family of the form
$q_{x}^{\star}(z)$ $=\left[1+(1-\rho)\left(\eta(z)-\ell(x;z)-A_{\rho}\right)\right]_{+}^{\frac{1}{\rho-1}}$ (17)
$=\exp_{2-\rho}\left[\eta(z)-\ell(x;z)-A_{\rho}\right],$ (18)
where $A_{\rho}$ is a normalization function and $a]_{+}=\max\{0,a\}$ is the positive part operator or rectified linear unit (ReLU). Notice that the ReLU operation may assign exactly zero probability (density) to some elements of $\mathcal{Z}$, and the support of $q_{x}^{\star}$ may depend on $x$, something that is impossible in any exponential family.
Figure 1: 1D logistic regression with a $\mathcal{N}(0,1)$ prior with $(2-\rho)$-Gaussian posteriors. Sparse posteriors (larger $\rho$) are closer to the MAP solution, as their truncation avoids sampling any points with high loss. In contrast, heavy-tailed posteriors move farther away from the high loss regions.
We stress that FYVI, with Tsallis $\rho$-negentropy regularization and $\rho>1$ leads to a qualitatively different type of sparsity from what is obtained by using sparsity-inducing priors or regularizers, such as LASSO (equivalently, Laplacian priors [LASSO, Hastie]), or spike-and-slab priors [spike]. Sparsity-inducing priors are used to encourage the estimated object—typically a vector of parameters of coefficients—to be sparse, but the associated posterior distributions have full support. For example, the Bayes posterior resulting from a Laplacian sparsity-inducing prior and a Gaussian observation model has full support. The reason why it is said that such priors induce sparsity is that the posterior mode (the maximum a posteriori or MAP estimate) can be exactly zero. In contrast, using any non-sparsity-inducing Gaussian prior and a Gaussian observation model in FYVI with, say, Tsallis $2$-negentropy regularization, yields a generalized posterior with bounded support.
3.2 Fenchel-Young Free Energy, Evidence, Evidence Lower Bound, and Posterior
By analogy with classical VI (4) the objective in (16) can be interpreted as a Fenchel-Young variational free energy (FYVFE). It is useful to include an additional parameter $\beta$ weighting the KLD regularizer in (5) and the FY discrepancy below, as in $\beta$-VAEs [Higgins2017betaVAELB]:
$F_{\ell,\Omega,\beta}(q,\eta;x)=\underset{z\sim q}{\mathbb{E}}\left[\ell(x;z)\right]+\beta D_{\Omega}(\eta;q).$ (19)
The parameter $\beta$ controls the strength with which the agreement between $q$ and $\eta$ is encouraged, and can also be seen from an information-theoretic perspective as controlling a rate-distortion trade-off [alemi2018information]. In the sequel, omitting $\beta$ means assuming $\beta=1$.
By analogy with (3), the minimum of 19 is the (negative) FY evidence,
$J_{\ell,\Omega}(x,\eta)=\min_{q\in\mathcal{Q}}F_{\ell,\Omega}(q,\eta;x)=F_{\ell,\Omega}(q_{x}^{\star},\eta;x),$
and the minimizer $q_{x}^{\star}$ is the FY posterior. By construction, the FY evidence, $J_{\ell,\Omega}(x,\eta)$, is the maximum achieved by $F_{\ell,\Omega}(q,\eta;x)$, for any $q\in\Delta(\mathcal{Z})$. Therefore the negative free energy $F_{\ell,\Omega}(q,\eta;x)$ is a FY evidence lower bound (FYELBO). For $\mathcal{Q}=\Delta(\mathcal{Z})$, $\Omega(q)=\Omega_{1}(q)=-H(q)$, and $\ell(x;z)=-\log p_{X|Z}(x|z)$, the FY evidence reduces to the regular log-evidence, $\log p_{X}(x)$, and the FY posterior recovers the standard Bayesian posterior, $q_{x}^{\star}=p_{Z|X}(\cdot|x)$.
In classical Bayesian analysis, a prior on the model space allows us to use the evidence for model comparison or hypothesis testing [Bayes_factors]. While the FY evidence may lack the usual statistical interpretations such as a marginal likelihood, Bayesian model selection, validation, and hypothesis testing, we still expect it to be useful in practical settings. For instance, it may work well in deep learning, where the usual marginal likelihood has been successfully applied despite a lack of Bayesian modeling [immer2021scalable].
3.3 Fenchel-Young Expectation Maximization (FYEM)
Here, we show for the example of the expectation maximization algorithm how our new FYVI framework leads to the natural generalization of existing algorithms to accommodate flexible posterior distributions. We call the resulting method FYEM. Let $x$ be an $N$-dimensional vector of observations $x=(x_{1},...,x_{N})\in\mathcal{X}^{N}$, and $z$ a vector of associated latent variables $z=(z_{1},...,z_{N})\in\mathcal{Z}^{N}$, with corresponding variational distributions $q=(q_{1},...,q_{N})\in\Delta(\mathcal{Z})^{N}$. Let $\ell_{\theta}(x;z)=\sum_{i=1}^{N}\mathcal{L}_{\theta}(x_{i};z_{i})$ be a parameterized loss function: the sum of sample-wise losses, where $\theta$ is a vector of parameters to be estimated. Here, we minimize the FYVFE (19) with respect to $q$ and $\theta$,
$\hat{q},\hat{\theta},\hat{\eta}=\arg\min_{q,\theta,\eta}\,\sum_{i=1}^{N}F_{\ell_{\theta},\Omega}(q_{i},\eta;x_{i})$ (20)
$=\arg\min_{q,\theta,\eta}\,\sum_{i=1}^{N}\Bigl(\,\underset{z_{i}\sim q_{i}}{\mathbb{E}}\left[\mathcal{L}_{\theta}(x_{i};z_{i})\right]+D_{\Omega}(\eta;q_{i})\Bigr),$ (21)
and also with respect to $\eta$, which is a score function.
Alternating minimization w.r.t. $q$ and $(\theta,\eta)$ yields an EM-style algorithm: In the E-step we compute the posteriors of the latent variables; as in standard EM, each $q_{i}$ can be obtained independently of the others. In the M-step we update the estimates of $\theta$ and $\eta$.
E-step:
Given the current $\hat{\theta}$ and $\hat{\eta}$, update $\hat{q}_{i}$ by minimizing (21),
$\hat{q}_{i}^{\mbox{new}}$ $=\arg\min_{q_{i}}\underset{z_{i}\sim q_{i}}{\mathbb{E}}\left[\mathcal{L}_{\hat{\theta}}(x_{i};z_{i})\right]+D_{\Omega}(\hat{\eta};q_{i})$
$=\arg\max_{q_{i}}\underset{z_{i}\sim q_{i}}{\mathbb{E}}\left[\hat{\eta}(z_{i})-\mathcal{L}_{\hat{\theta}}(x_{i};z_{i})\right]-\Omega(q_{i})$
$=\arg\min_{q_{i}}D_{\Omega}(\hat{\eta}-\mathcal{L}_{\hat{\theta}}(x_{i};\cdot);q)$
$=\Pi_{\Omega}(\hat{\eta}-\mathcal{L}_{\hat{\theta}}(x_{i};\cdot)),$ (22)
for i=1,,Ni=1,...,N, where $\Pi_{\Omega}$ is the $\Omega$-regularized prediction map (11).
Figure 2: We compare qualitatively the standard, sparse, and hard versions of EM for GMM-based clustering. Data points are colored by their true labels, with x indicate noise points. Colored ellipses denote the data generating covariances; they are the same across the three panels. Black ellipses are level curves of the estimated GMM components and they illustrate the effect of E-step sparsity on the fit. Our sparse EM algorithm combines the best of both worlds: It is robust to the outliers while still retaining soft-assignments as in the original EM.
M-Step:
Given $\hat{q}_{i}^{\mbox{new}}$, update $\hat{\theta}$ and $\hat{\eta}$. To update $\hat{\theta}$, we solve
$\hat{\theta}^{\mbox{new}}=\arg\min_{\theta}\sum_{i=1}^{N}\underset{z_{i}\sim\hat{q}_{i}^{\mbox{new}}}{\mathbb{E}}[\mathcal{L}_{\hat{\theta}}(x_{i};z_{i})],$
which depends on the particular form of $\mathcal{L}_{\hat{\theta}}$. To update $\hat{\eta}$, we recall from Eq. 7 that $D_{\Omega}(\eta;q)=\Omega^{*}(\eta)-\mathbb{E}_{z\sim q}[\eta(z)]+\Omega(q)$, and thus
$\hat{\eta}^{\mbox{new}}$ $=\arg\min_{\eta}\,\sum_{i=1}^{N}D_{\Omega}\bigl(\eta;\hat{q}_{i}^{\mbox{new}}\bigr)$
$=\arg\min_{\eta}\,\Omega^{*}(\eta)-\frac{1}{N}\sum_{i=1}^{N}\underset{z_{i}\sim\hat{q}_{i}^{\mbox{new}}}{\mathbb{E}}[\eta(z_{i})]$
$=\arg\max_{\eta}\,\underset{z\sim\bar{q}}{\mathbb{E}}[\eta(z)]-\Omega^{*}(\eta),$ (23)
where $\bar{q}=(1/N)\sum_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}$ by linearity of the expectation.
Assuming $\mathcal{Z}=\{1,...,K\}$ is a discrete set, $\eta$ is simply a vector $\eta=(\eta_{1},...,\eta_{K})\in\mathbb{R}^{K}$ and $q_{i}=(q_{i_{1}},...,q_{i_{K}})\in\Delta(\mathcal{Z})$ is a probability distribution. In this case, $\mathbb{E}_{z\sim\bar{q}}[\eta(z)]=\langle\bar{q},\eta\rangle=\sum_{k=1}^{K}\bar{q}_{k}\eta_{k}$, thus
$\hat{\eta}^{\mbox{new}}\in\arg\min_{\eta}\Omega^{*}(\eta)-\langle\bar{q},\eta\rangle,$ (24)
where we have written the update as an inclusion because, as shown by martins2022sparse, the minimizer may not be unique and is achieved by any $\hat{\eta}$ satisfying $\Pi_{\Omega}(\hat{\eta})=\bar{q}$. If $\Omega$ is the Shannon negentropy, then $\Pi_{\Omega}$ will be the softmax transformation, and $\hat{\eta}^{\mbox{new}}=(\log\bar{q}_{1},...,\log\bar{q}_{K})+C$, where $C$ is an arbitrary constant, which without loss of generality we can set to zero since it does not affect the algorithm. Under this choice, the FYEM recovers a regular EM algorithm. If $\Omega$ is the Tsallis 2-negentropy $\Omega_{2}$, then $\Pi_{\Omega}(\eta)$ is the Euclidean projection of $\eta$ onto the simplex, and any $\hat{\eta}$ of the form $\hat{\eta}^{\mbox{new}}=(\bar{q}_{1},...,\bar{q}_{K})+C$, where $C$ is again an arbitrary constant that we set to zero, is a solution of (24) (although there may exist others). If $\Omega=0$, we show in Appendix B that the above FYEM recovers the classical hard EM algorithm.
3.4 Gradient Backpropagation
We also propose a simple and versatile gradient backpropagation method to perform FYVI for general models. We use it in two of the experiments below. There, we optimize the FYELBO using amortized inference [Gershman], where the variational distribution is parametrized by an encoder network that takes $x$ as input. We denote the resulting variational distribution by $q_{\phi}(z|x)$, where $\phi$ are the inference network parameters. Instead of minimizing (21) through alternating minimization, we use end-to-end gradient backpropagation combined with a reparametrization trick to update the encoder parameters $\phi$, the decoder parameters $\theta$, and optionally any parameters associated with $\eta$ (our FYVAEs use standard fixed priors). It is also obviously possible to use a score function estimator instead of reparametrization. We also note that generalized natural-gradient algorithms could be employed here.
4 EXPERIMENTS
We illustrate FYVI through several experiments. The first is GMM-based clustering, where we recover the standard and hard versions of EM with particular choices of entropic regularizer, in addition to a new sparse EM algorithm. [^2] This sparse variant of EM is a different instance of FYVI than those proposed by dukkipati2016mixture and inoue2013q, which are standard mixtures of $(2-\rho)$-Gaussians. We also use FYVI to train $\beta$-VAEs with compact support posteriors on MNIST and FashionMNIST with both sparse and standard multivariate Bernoulli observation models. Finally, we combine a Gaussian prior with sparse observation models to model a sparse empirical distribution in a neural variational document model.
Our focus is not computational scale. Computational cost depends on the particular model, entropic index $\rho$, and parametrization. Therefore, the computational cost depends on model-specific choices, rather than the generalization of standard VI to FYVI. Tradeoffs between FY and KL regularization similarly depend on model specification; there is no general comparison to be made. In our experiments, computational tradeoffs between FY regularizers and KL regularizers are negligible. There are efficient algorithms for computing sparsemax ($\rho=2$) and $1.5$-entmax; otherwise computing softmax is faster. We use reparametrization tricks in all the VAE experiments: The classical Gaussian one [Blei2017], and a $(2-\rho)$-Gaussian reparametrization [martins2022sparse]. We keep computational resources equal between compared models.
Figure 3: We vary $\rho\in\{0.1,0.5,0.9,1.0,1.1,1.5,2.0,3.0\}$ and plot from left to right the resulting adjusted mutual information, adjusted Rand index, and silhouette score. In the right-most panel we quantify the E-step sparsity by counting the number of clusters assigned 0 mixing proportion at the last iteration of each EM run and averaging across samples. Since the total number of clusters is four, the greatest number of components assigned 0 proportion is three. The shaded regions represent standard errors from the mean values across 5 seeds.
4.1 FYEM for GMMs
We set $\mathcal{L}_{\theta}(x_{i},z_{i})=-\log\mathcal{N}(x_{i};\mu_{z_{i}},\Sigma_{z_{i}})$ as Gaussian negative log-likelihoods. Choosing $\Omega=\Omega_{1}$ recovers the standard EM algorithm for GMM estimation, where all the components of $\hat{q}_{i}^{\mbox{new}}$ are strictly positive. We also use other regularizers, $\Omega_{\rho}$ with $\rho>1$, which yield a new adaptively sparse EM algorithm for GMMs. In Appendix B we review the classical EM and hard EM algorithms compared in Figure 2, and give details of how the clusters are generated and evaluated in Table 1. The sparse EM algorithm interpolates between the well known standard and hard EM variants and is thereby robust to the outliers, similarly to hard EM, while still retaining soft-assignments as in the classical EM. We do not report the log marginal likelihood since this is maximized by standard EM, while sparse and hard EM optimize the FYVI objectives. Figure 3 illustrates how average sparsity and the unsupervised clustering evaluation metrics we use vary with $\rho$.
Setting $\Omega=\Omega_{\rho}$ to the Tsallis negentropy (14) with $\rho>1$, yields the new sparse version of (22):
${\hat{q}_{i}^{\mbox{new}}}=$ $\rho\mbox{-entmax}\Bigl(\frac{\pi_{1}^{\rho-1}}{\rho-1}+\log\bigl(\mathcal{N}(x_{i};\hat{\mu}_{1},\hat{\Sigma}_{1})\bigr),$...,
$\frac{\pi_{K}^{\rho-1}}{\rho-1}+\log\bigl(\mathcal{N}(x_{i};\hat{\mu}_{K},\hat{\Sigma}_{K}\bigr)\Bigr);$ (25)
see [peters2019sparse] for efficient implementations of the $\rho\mbox{-entmax}$ function for certain values of $\rho$.
When $\rho>1$ the output of $\rho\mbox{-entmax}$ can have both zero and non-zero components. The novel feature of sparse EM is the possible presence of zeros in $\hat{q}_{i}$, meaning that sample $x_{i}$ will not contribute to the update of the parameter estimates of the corresponding components. In standard EM all samples affect the parameter estimates of all the components. In hard EM, each sample only affects the parameters of the “closest” component, as in the case of $K$-means clustering (with the important difference that $K$-means clustering does not estimate component covariances).
Table 1: Clustering quality metrics (higher is better) for GMM-based clustering using the standard, hard, and sparse (ours) versions of EM; $\rho=2$ in sparse EM. Priors are learned according to Eq.(23). Values are averaged over 5 random seeds $\pm$ standard errors.
E-Step Adj. MI Adj. RI Silhouette
Classic .606 $\pm$ .005 .531 $\pm$ .002 .345 $\pm$ .014
Hard .537 $\pm$ .039 .348 $\pm$ .019 .207 $\pm$ .057
Sparse .636 $\pm$ .008 .476 $\pm$ .017 .393 $\pm$ .018
Table 2: VAE $\ell_{1}$ reconstruction error on MNIST and FashionMNIST datasets over 5 seeds $\pm$ std. error.
Observation Model Latent Variable Mnist FashionMnist
Product of independent Bernoulli’s, $\rho^{\prime}=1$ Gaussian, $\rho=1$ 13.236 $\pm$ .004 34.049 $\pm$ .019
Biweight, $\rho=1.5$ 12.065 $\pm$ .021 32.480 $\pm$ .018
Epanechnikov, $\rho=2$ 12.086 $\pm$ .012 32.537 $\pm$ .025
Product of $\Psi_{2}$ sparse binary distributions, $\rho^{\prime}=2$ Gaussian, $\rho=1$ 12.567 $\pm$ .017 32.881 $\pm$ .027
Biweight, $\rho=1.5$ 9.261 $\pm$ .009 25.970 $\pm$ .086
Epanechnikov, $\rho=2$ 9.144 $\pm$ .020 25.851 $\pm$ .072
4.2 Fenchel-Young $\beta$-VAE
In Kingma2014’s classic VAE a Gaussian variational posterior and a $\mathcal{N}(0,I)$ prior are combined with a product of independent Bernoulli’s as observation model, representing the pixels of MNIST digit images. The standard $\beta$-VAE baseline objective is reviewed in in Appendix C. We experiment with compact support $(2-\rho)$-Gaussian333martins2022sparse call this family $\beta$-Gaussian. variational posteriors. We also compare the baseline product of Bernoulli’s decoder with a product of binary $\Psi_{2}$ deformed exponential family distributions. The FYELBO objective is given as follows:
$\sum\limits_{i=1}^{N}\underset{z\sim q_{i}}{\mathbb{E}}\left[\ell_{\Psi_{\rho^{\prime}}}(\theta(z);x_{i})\right]+\beta D_{\Omega_{\rho}}(\eta_{\mathcal{N}(0,\textnormal{I})};q_{i}).$
[…]
It is interesting to note that sparsity in the encoder and in the decoder separately contribute to improved performance. We report average $\ell_{1}$ reconstruction error as a means of comparing the different models, since the FYELBO values themselves are not comparable. In all cases we tune $\beta$ by cross-validation to prevent the posterior from collapsing to the prior.
In Appendix D we report improved $\ell_{1}$ reconstruction error in a small document model VAE, using sparsity only in the observation model. This can be seen as a generalized Bayesian treatment of sparse observation models, falling under the cases of generalized VI described in [bissiri2016general, JMLR:v24:22-0291, knoblauch2022optimization]. In preliminary experiments with sparse posteriors we did not observe improvements over the baseline NVDM of miao2016neural. This is intuitively reasonable: While the empirical distribution over words is sparse, one latent Multivariate Gaussian associated with each Multinomial (or $\Omega_{\rho}$-sparse) observation through amortized variational inference does not lead to improved reconstruction if its domain is truncated. In the image VAE data, similar patterns of pixels are 0 across images, for example on the boarders, while in the document model VAEs there is no word of the vocabulary that ought to be assigned no probability mass in all “topic dimensions”.
5 CONCLUSION
We have introduced Fenchel-Young variational inference (FYVI), expanding the class of variational posteriors to the $(2-\rho)$-deformed exponential family. FYVI exploits a general class of entropic regularizers to obtain sparse discrete and continuous posterior distributions, and distinctly extends classical Bayesian inference compared to previous generalized Bayesian and variational inference proposals like those of bissiri2016general, knoblauch2022optimization, and daudel2023monotonic. We focus on cases where the entropic index $\rho$ is greater than $1$, which leads to sparse posteriors previously unexplored in the related literature. Note that the nature of sparsity of $(2-\rho)$-exponential family posteriors is qualitatively distinct from the sparse MAP estimates that are obtained through sparsity-inducing priors. We define and lower bound the FY variational free energy (FYVFE), and describe EM-style and amortized VI algorithms for estimating arbitrary generative FYVI models. We derive and experiment with a sparse EM algorithm for Gaussian mixtures, demonstrating better clustering quality than classical and hard versions of EM on synthetic data with overlapping clusters and outliers. In our image FYVAE experiments with continuous finite support posteriors, we obtain the best performance when combining a latent $(2-\rho)$-Gaussian posterior with a sparse support observation model. We also show how to treat sparse observation models from a generalized Bayesian perspective.
FYVI can be directly extended to structured latent variables by using SparseMAP [niculae2018sparsemap] posteriors; intuitively these are sparse analogues to conditional random fields. In future work we plan to extend our analysis and experiments to the non-trivial case of heavy tailed posteriors ($\rho$ < 1). Further, FYVI is a necessary step toward characterizing conjugate FY families and generalized natural gradient descent, to account for the geometry of the space of deformed exponential family distributions. Another interesting direction is to connect FYVI with variational Rényi objectives [daudel2023alpha, daudel2023monotonic]. The recently popular $\lambda$-deformation [zhang2022lambda, wong2022tsallis, guilmeau2023variational], which aims to treat Rényi and Tsallis entropies in a unified way, may provide an illuminating bridge among alternative generalized VI perspectives. Finally, connections to the recent works from online learning by alquier2021non and PAC-Bayes [alquier2016properties, alquier2024user] are also an interesting prospect.
Appendix
Appendix A RELATION BETWEEN FY EXPRESSIONS, $\alpha$- AND $q$-DIVERGENCES
In this section, we discuss how the Fenchel-Young discrepancies introduced in the main paper are related to some well-known divergences between probability distributions from the literature. This not an exhaustive review, rather we give some context about closely related and easily confused quantities. We begin by introducing the deformed $\log$ and its inverse as these functions are necessary for defining the Tsallis relative entropy, which is closely related to amari2011geometry’s $\alpha$-divergence.
$\rho$-deformation.
The Tsallis entropy is based on the $\rho$-deformed logarithmic and exponential functions for an entropic index $\rho\in\mathbb{R}\setminus\{1\}$ — this index is called $q$ by tsallis2009introduction, amari2011geometry, naudts2009q, and $\alpha$ by martins2022sparse:
$\log_{\rho}(x):=\frac{x^{1-\rho}-1}{1-\rho},\quad x>0,\quad\exp_{\rho}(x):=\left[1+(1-\rho)x\right]_{+}^{\frac{1}{1-\rho}},\quad x\in\mathbb{R}$ (26)
where $\cdot]_{+}:=\max\{\cdot,0\}$. In the limit $\rho\rightarrow 1$ these converge to $\log(x)$ and $\exp(x)$. Note that multiplication and division inside deformed logarithms is modified as follows:
$\log_{\rho}\left(\frac{x}{y}\right)=\frac{\log_{\rho}x-\log_{\rho}y}{1+(1-\rho)\log_{\rho}y},\quad\log_{\rho}(xy)=\log_{\rho}x+\log_{\rho}y-(1-\rho)\log_{\rho}x\log_{\rho}y.$ (27)
$D_{\Omega}$ and related divergences.
Eq. 14 in the main paper defines the Tsallis $\rho$-negentropy scaled by $\frac{1}{\rho}$ consistently with martins2022sparse: $\Omega_{\rho}(q)=-\frac{1}{\rho}S_{\rho}(q)$ where $S_{\rho}(q)=\frac{1}{1-\rho}\mathbb{E}_{q(z)}\left[q(z)^{\rho-1}-1\right]$ is the Tsallis entropy. Additionally, we note the alternative forms, using the fact that $\log_{\rho}\left(\frac{1}{q}\right)+\log_{2-\rho}q=0$,
$S_{\rho}(q)=\frac{1}{1-\rho}\mathbb{E}_{q(z)}[q(z)^{\rho-1}-1]=\mathbb{E}_{q(z)}\log_{\rho}\frac{1}{q(z)}=-\mathbb{E}_{q(z)}\log_{2-\rho}q(z).$ (28)
The Tsallis $\rho$-relative entropy, given by Eq. 39 in amari2011geometry is a scalar multiple of their $\alpha$-divergence (equivalent for $\alpha=1-2\rho$):
${D}_{\rho}^{\textnormal{Tsallis}}(q,p)$ $=-\underset{z\sim q(z)}{\mathbb{E}}\log_{\rho}\frac{p(z)}{q(z)}=-\int q(z)^{\rho}\left[\log_{\rho}p(z)-\log_{\rho}q(z)\right]dz$ (29)
$=\frac{1}{1-\rho}\left[1-\int q(z)^{\rho}p(z)^{1-\rho}dz\right].$ (30)
amari2011geometry use a $\rho$-divergence (see their Eq. 40), which is a Bregman divergence based on the convex $\rho$-cumulant function $A_{\rho}$, between two $\rho$-exponential densities with form $\log_{\rho}q=\eta_{q}^{\top}z-A_{\rho}(\eta_{q})$, where density $q$ is parametrized by $\eta$ and we index canonical parameter $\eta$ by the density to write the canonical $\rho$ divergences as a Bregman divergence below. Differentiating $A_{\rho}$ yields a dual parametrization of the $\rho$-exponential family that is equal to the escort mean $q=\mathbb{E}_{\tilde{q}}[z]$, where $\tilde{q}(z;\eta)=\frac{q(z;\eta)^{\rho}}{\int q(z;\eta)^{\rho}dz}$ is the escort distribution and the escort expectation parameter $q$ is denoted by the name of the distribution. amari2011geometry show that the $\rho$-divergence is a “conformal transformation” of ${D}_{\rho}^{\textnormal{Tsallis}}$. Note also that if we multiply Amari’s $\alpha$-divergence [amari2016information] by $\frac{\rho}{\int p_{i}^{\rho}}$ (for $\rho=\frac{1-\alpha}{2}$) we retrieve the $\rho$-divergence:
$\begin{split}D_{\rho}(q,p)&=\frac{\rho}{\int q_{i}^{\rho}}\underbrace{\frac{1}{\rho}D^{\textnormal{Tsallis}}_{\rho}(q,p)}_{\textnormal{Amari's }\alpha\textnormal{-divergence}}=\frac{-\int q_{i}(\log_{\rho}p_{i}-\log_{\rho}q_{i})}{\int q_{i}^{\rho}}=\underset{{\tilde{q}}}{\mathbb{E}}[\log_{\rho}q_{i}-\log_{\rho}p_{i}]\\ &=\underset{{\tilde{q}}}{\mathbb{E}}[\eta_{q}^{\top}z-A_{\rho}(\eta_{q})-(\eta_{p}^{\top}z-A_{\rho}(\eta_{p}))]=\eta_{q}^{\top}q-A_{\rho}(\eta_{q})-(\eta_{p}^{\top}q-A_{\rho}(\eta_{p}))\\ &=A_{\rho}(\eta_{p})-A_{\rho}(\eta_{q})-\underbrace{q^{\top}}_{=\nabla A_{\rho}(\eta_{q})}(\eta_{p}-\eta_{q})=\textnormal{Bregman}_{{A_{\rho}}}(\eta_{p},\eta_{q})\\ &=A^{*}_{\rho}(q)+A_{\rho}(\eta_{p})-\underset{{\tilde{q}}}{\mathbb{E}}[z^{\top}\eta_{p}].\end{split}$ (31)
For the expressions in the last line we use the Legendre duality $\eta_{q}^{\top}q-A_{\rho}(\eta_{q})=A^{*}_{\rho}(q)$.
Maximum Tsallis $\rho$-Entropy.
There are two alternative approaches to derive the same $\rho$-exponential family: Maximize the $(2-\rho)$-entropy subject to linear expectation constraints, or maximize the $\rho$-entropy subject to escort expectation constraints. amari2011geometry follow the latter approach and are interested in characterizing the divergence between densities $q,p$ that belong to the $\rho$-expfam obtained by maximizing the Tsallis $\rho$-entropy subject to escort distribution constraints. In this paper, we focus on the $(2-\rho)$-exponential family of distributions that arises from the Tsallis $\rho$-MaxEnt principle subject to standard linear expectation constraints, following naudts2009q and martins2022sparse. These densities are the $\Omega_{\rho}$-regularized prediction maps $p(z;\eta)=\exp_{2-\rho}[\eta^{\top}z-A_{\rho}(\eta)]$ ([martins2022sparse, Def. 2]). Overloading $\eta$ to denote both a function and canonical parameters, the scoring function is $\eta(z)=\eta^{\top}z$. The FY expression in Eq. (15) above, induced by the negentropy $\Omega$ compares two members of the $(2-\rho)$-exponential family: $D_{\Omega}(\eta;q)=\Omega_{\rho}(q)-\Omega_{\rho}(p(z;\eta))-\eta^{\top}(\mathbb{E}_{p_{\eta}}[z]-\mathbb{E}_{q}[z])$ ([martins2022sparse, Eq. 14]).
Starting with regularized prediction maps like [martins2022sparse, Def. 2] the index $\rho$ of the Tsallis entropy is specified first, and then the associated deformed family is defined using $\exp_{2-\rho}$, based on the $\rho$-MaxEnt problem under standard constraints [martins2022sparse, Eq. 10]. On the other hand, if the index $\rho$ is used to write a $\rho$-exponential family first, then the associated Tsallis entropy that is maximized, again with standard constraints, is the $(2-\rho)$-entropy (see Chapters 7 and 8 in naudts2011generalised and wada2005connections, wada2012relationships).
Proposition 1 (Fenchel-Young VI and Max Tsallis Entropy Equivalence).
Let $\mathcal{Q}$ be a set of probability distributions over a space $\mathcal{Z}$, let $\ell(x;z)$ be a measurable loss function, and let $\eta:\mathcal{Z}\rightarrow\mathbb{R}$ be a scoring function (possibly a log-prior). $\Omega:\mathcal{Q}\to\mathbb{R}\cup\{+\infty\}$ be a proper, lower semi-continuous, convex regularizer that induces a Fenchel-Young discrepancy:
$D_{\Omega}(\eta;q):=\Omega(q)+\Omega^{*}(\eta)-\mathbb{E}_{z\sim q}[\eta(z)].$
Then the variational problem
$q_{x}^{\star}=\arg\min_{q\in\mathcal{Q}}\penalty 10000\ \mathbb{E}_{z\sim q}[\ell(x;z)]+L_{\Omega}(\eta;q)$
is equivalent to a generalized Tsallis maximum entropy problem under standard expectation constraints, where the regularizer $\Omega$ is chosen as the negative Tsallis $\rho$-entropy for $\rho>1$:
$\Omega_{\rho}(q)=-\frac{1}{\rho}S_{\rho}(q)=\frac{1}{\rho}\frac{1}{\rho-1}\left(1-\int q(z)^{\rho}dz\right).$
Moreover, the solution $q_{x}^{\star}$ belongs to the $(2-\rho)$-exponential family, i.e.,
$q_{x}^{\star}(z)$ $=\left[1+(1-\rho)\left(\eta(z)-\ell(x;z)-A_{\rho}\right)\right]_{+}^{\frac{1}{\rho-1}}$
$=\exp_{2-\rho}\left[1+\eta(z)-\ell(x;z)-A_{\rho}\right],$
where $A_{\rho}$ is a normalization function.
Proof.
The variational problem
$q_{x}^{\star}=\arg\min_{q\in\mathcal{Q}}\mathbb{E}_{z\sim q}\left[\ell(x;z)\right]+D_{\Omega}(\eta;q)$
defines the $\Omega$-regularized prediction map with scoring function $\eta(z)-\ell(x;z)$
$q_{x}^{\star}(z)=\Pi_{\Omega}(\eta(z)-\ell(x;z))=\arg\max_{q\in\mathcal{Q}}\mathbb{E}_{z\sim q}\left[\eta(z)-\ell(x;z)\right]-\Omega(q)$
that maximizes the $\rho$-Tsallis entropy subject to an expectation constraint [martins2022sparse, App. C.4]. Using [martins2022sparse, Prop. 8] we can express $q_{x}^{\star}(z)$ as a member of the $(2-\rho)$-exponential family. ∎
Appendix B CLUSTERING WITH FY EM
Standard EM.
The FY EM algorithm described in $\S$3.3 recovers standard EM for GMMs when: $x_{1},...,x_{N}\in\mathbb{R}^{d}$ are $d$-variate observations and the latent variables $z_{1},...z_{N}\in\mathcal{Z}=\{1,...,K\}$ indicate to which component each observation $x_{i}$ belongs; $q_{i}=\bigl(q_{i}(1),...,q_{i}(K)\bigr)\in\!\Delta(\mathcal{Z})$ is the variational approximation of the distribution of $z_{i}$; $\eta\in\mathbb{R}^{K}$ contains the logarithms of the prior probabilities of the $K$ components, denoted $\pi_{1},...,\pi_{K}$, i.e., $\eta_{1}=\log\pi_{1},...,\eta_{K}=\log\pi_{K}$; $\ell(x_{i},z,\theta)=-\log\mathcal{N}(x_{i};\mu_{z},\Sigma_{z})$, for z=1,,Kz=1,...,K, where $\mu_{z}$ and $\Sigma_{z}$ are, respectively, the mean and covariance of component $z$; and $\Omega(q)=-H(q)$, thus $\nabla\Omega^{*}=$ softmax.
The E-step in (22) then takes the familiar form
$\hat{q}_{i}^{\mbox{new}}(z)=\mbox{softmax}\bigl(\eta_{z}+\log\mathcal{N}(x_{i};\hat{\mu}_{z},\hat{\Sigma}_{z})\bigr)=\pi_{z}\;\mathcal{N}(x_{i};\hat{\mu}_{z},\hat{\Sigma}_{z})\biggl(\sum_{j=1}^{K}\pi_{j}\;\mathcal{N}(x_{i};\hat{\mu}_{j},\hat{\Sigma}_{j})\biggr)^{-1}\!,$
for i=1,,N,i=1,...,N, z=1,,Kz=1,...,K, where $\hat{\mu}_{z}$, $\hat{\Sigma}_{z}$ are the current parameter estimates. This is the E-step of standard EM for GMM estimation [Figueiredo2002]. The M-step corresponds to maximum weighted log-likelihood estimation, yielding, for $z_{i}=1,...,K$,
$\hat{\mu}_{z}^{\mbox{new}}$ $=\frac{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}(z)\;x_{i}}{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}(z)},\quad\hat{\Sigma}_{z}^{\mbox{new}}=\frac{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}(z)\;(x_{i}-\hat{\mu}_{z}^{\mbox{new}})(x_{i}-\hat{\mu}_{z}^{\mbox{new}})^{\top}}{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}(z)},\hat{\eta}_{z}^{\mbox{new}}$ $=\log\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}(z)-\log N.$
Hard EM.
Setting $\Omega(q)=0$, i.e., removing the regularizer, yields a different E-step:
$\hat{q}_{i}^{\mbox{new}}\!=\mbox{argmax}\bigl(\pi_{1}\;\mathcal{N}(x_{i};\hat{\mu}_{1},\hat{\Sigma}_{1}),...,\pi_{K}\;\mathcal{N}(x_{i};\hat{\mu}_{K},\hat{\Sigma}_{K})\bigr),$
where $\mbox{argmax}(u_{1},...,u_{K})$ of a tuple of $K$ numbers returns a vector with $1/m$ in the entries with the same indices as the $m$ maxima and zero everywhere else. If the maximum is unique, this is just the one-hot representation of the maximizer. This hard version of EM coincides with the classification EM (CEM) algorithm proposed by Celeux. The M-step is the same as in standard EM.
Experiment Details.
We generate $1000$ samples uniformly from four bivariate Gaussians with means at [1,1][-1,-1], $[0,0]$, $[1,1]$, and [1,1][1,-1]. The components overlap significantly, with covariance matrices 0.11I0.11\,I, 0.5I0.5\,I, 0.7I0.7\,I, and 0.9I0.9\,I. Additionally, we sample $100$ uniformly distributed random samples in $3,3]^{2}$ to simulate the presence of outliers, yielding a total of $1100$ data points. We compare the three EM variants described —standard EM, hard EM, and sparse EM— on unsupervised clustering of this dataset. The number of components in each model is set to $4$ and the initial parameters (means and covariances) are randomly initialized from a uniform distribution in [0, 0.1][0,\,0.1]. Each algorithm runs $200$ iterations, which we verified is more than enough to satisfy a tight convergence criterion.
We evaluate the quality of unsupervised clustering of each algorithm using three metrics: adjusted mutual information [Vinh], adjusted Rand index [Rand], and silhouette score [Rousseeuw]. Each model is trained and evaluated across five different random seeds, and the average and standard deviation of the metrics are reported in Table 1, showing that sparse EM outperforms standard and hard EM in two of three standard metrics. See also Figure 2 for a qualitative illustration. In Figure 3 we vary $\rho=0.1,0.5,0.9,1.0,1.1,1.5,2.0,3.0$ and evaluate the same three metrics when clustering with the corresponding EM. In the right most image of Figure 3 we verify that the sparsity of the E-step corresponds to the value of $\rho$ for the same values of $\rho$. We quantify the E-step sparsity by counting the number of clusters assigned zero mixing proportion at the last iteration of each EM run and averaging across samples. Since the total number of clusters is four, the greatest number of components assigned zero proportion is three.
Appendix C MNIST AND FMNIST FY VAES
The objective maximized by standard amortized VI w.r.t. $\theta$ (decoder parameters) and $\phi$ (encoder parameters), for a set of $N$ images is
$\sum\limits_{i=1}^{N}\underset{z\sim q_{i}}{\mathbb{E}}\bigl[\log p_{\theta}(x_{i}|z)\bigr]-\beta D_{\text{KL}}(q_{i}\|p_{Z}),$ (32)
where $q_{i}=\mathcal{N}\bigl(\mu_{\phi}(x_{i}),\mbox{diag}\bigl(\sigma^{2}_{\phi}(x_{i})\bigr)\bigr)$ and $p_{\theta}(x_{i}|z_{i})\sim\textnormal{Bernoulli}(\sigma(f_{\theta}(z_{i}))$; $\sigma$ denotes the sigmoid function, and $f_{\theta}$ a neurally parametrized function.
Experiment Details.
[…]
We report the $\ell_{1}$ reconstruction error in Table 2, i.e. the absolute distance between the model’s reconstructed samples and the original input images, evaluated on the test split for both MNIST and FashionMNIST datasets. The reconstruction error is the only standard and interpretable metric that is comparable across the VAE alternatives we consider. The FY ELBO has a log scale only in the standard VAE case, which is lost when either the observation or latent variable model is based on a Tsallis entropy with $\rho>1$. The first row of Table 2 corresponds to the standard VAE with a Gaussian latent variable. The Biweight and Epanechnikov densities are limited support distributions of the latent variable, with Tsallis entropies corresponding to $\rho=3/2$ $\rho=2$ respectively. The best reconstruction errors are obtained with latent $(2-\rho)$-Gaussian encoders combined with 2-entmax observation models.
Appendix D DOCUMENT FY $\beta$-VAE
Similarly with one of the cases we compared in the image VAE, namely the sparsemax observation model combined with a Gaussian latent variable in the previous section, we experiment with a FY $\beta$-VAE applied to unsupervised word embeddings to directly model the sparsity of the empirical distribution. In this experiment we focus on tuning the level of sparsity in the observation model rather than using only $\rho^{\prime}=2$. The baseline neural variational document model (NVDM) proposed by miao2016neural is a standard VAE (with prior $p_{Z}=\mathcal{N}(0,I)$): it models the joint probability of observed word counts (or proportions) in a document $x_{i}$, i=1,,Ni=1,...,N, and of continuous latent semantic embeddings $z_{i}$, by optimizing the same objective as (32) and replacing the multivariate Bernoulli with a Multinomial observation model, $p_{\theta}(x_{i}|z_{i})\sim\textnormal{Multinomial}(\textnormal{softmax}(Wz_{i}))$, where $\theta=W\in\mathbb{R}^{|D|\times K}$, $K$ is the dimension of the latent multivariate Gaussian, and $|D|$ is the size of the observed vocabulary. The model parameters are learned by stochastic gradient backpropagation.
Our FY $\beta$-VAE model uses the same multivariate Gaussian latent $q$ and prior $p_{Z}$, combined with a FY observation model $D_{\Psi}(\theta(z);x)$: $J(\phi,\theta)=\sum\limits_{i=1}^{N}\underset{Z\sim q}{\mathbb{E}}\left[D_{\Psi}(\theta(Z);x_{i})\right]+\beta D_{\text{KL}}(q_{i}\|p_{Z})$, where $\Psi$ is a Tsallis $\rho^{\prime}$-negentropy, dom($\Psi)=\Delta_{|D|-1}$, and $\rho^{\prime}$-entmax replaces softmax in the $D_{\Psi}$ decoder. Unlike the standard VAE formulation, the deformed exponential family distribution parametrized by $\theta(z)$ can assign zero probability to vocabulary categories that do not occur in a given document, directly modeling the sparse empirical distribution.
Experiment Details. We use the 20NewsGroups [^4] http://qwone.com/~jason/20Newsgroups/, [lang1995newsweeder]. dataset, which consists of 11,314 training samples and 7,531 test samples (documents, represented as bags of words). We assume a decoder parameterizing a sparse discrete distribution over the vocabulary generates all the words in a document independently, given a continuous Gaussian latent variable $Z$, i.e. a dense, continuous document representation. Following miao2016neural, we assume that each dimension in the latent space represents a “topic”, which is a simplification of the classical discrete latent variable topic model of [blei2003latent].
The first 1,000 test samples were used as the validation set for hyperparameter tuning. We perform a search over entmax $\rho^{\prime}$, which controls the decoder’s sparsity, learning rate, batch size, number of epochs, and the number of topics using Optuna [optuna2019] across 100 trials for both the baseline and FY NVDM models. Finally, we rerun the experiment with the best hyperparameters using five different random seeds and report the reconstruction error (RE) defined by the $\ell_{1}$ distance between the empirical data distribution and the learned observation model distribution. For the baseline NVDM model over 100 Optuna trials the hyperparameters that yield the best reconstruction error of 1.484 reported in Table 3 are batch size = 64, number of alternating epochs updating the decoder and encoder parameters = 1. For the Fenchel-Young $\beta$-VAE over 100 Optuna trials the hyperparameters that yield the best reconstruction error of 1.384 reported in Table 3 are $\rho^{\prime}=1.314$, batch size = 1024, number of alternating epochs updating the decoder and encoder parameters = 3. In both cases $\beta=0.01$, learning rate $5\times 10^{-5}$, number of training epochs = 500, and dimension of the latent Gaussian = 50. Similarly to the image VAE experiments we use $\beta=0.01$ ensures that the regularization term is non trivially small. Both models converge based on training objective curves, and in preliminary experiments we empirically verified that for $\rho^{\prime}$ close to 1, training and validation behavior was similar to the NVDM baseline. Again the objectives are are not numerically comparable to each other, nor is expected $\Psi_{r}^{\prime}$ term with the log scale regularization term. Again $\ell_{1}$ reconstruction error is a metric that is comparable between the NVDM and the FY NVDM.
Table 3: Average reconstruction error over 5 seeds for each model normalized by the number of examples in the test set (6505).
Model Average RE
NVDM 1.4984 ± .0004
FY VAE 1.3384 ± .0004
Our sparse support observation model FYVAE outperforms the NVDM baseline method [miao2016neural] by 10.68% in absolute reconstruction error, as reported in Table 3. Figure 4 shows average coherence scores, measured by the pairwise similarity between embedding representations of the top-$k$ words for each latent “topic” dimension, using a sentence-transformer model [reimers-gurevych-2019-sentence]. [^5] We use the all-mpnet-base-v2 model, which is available at huggingface.co/sentence-transformers/all-mpnet-base-v2, under Apache-2.0. FYVAE consistently achieves better or similar coherence scores across varying values of $k$ (number of words) than the NVDM model, confirming that the latent document representations learned by FYVAE are more semantically coherent and better aligned with meaningful word groupings.
Figure 4: Topic coherence scores for the standard NVDM and the FY VAE models.